### v1

# Weighted Covariance Intersection for Range-based Distributed Cooperative Localization of Multi-Agent Systems

Abstract
Precise localization of multi-agent systems (MAS) in harsh environments is a critical challenge for swarm applications, and cooperative localization is considered a key solution to this issue. Among all solutions, distributed cooperative localization (DCL) has garnered widespread attention due to its robustness and scalability. The main challenge of DCL lies in how to fuse relative measurements between agents under unknown correlations. To address this, covariance intersection (CI) was introduced to DCL. However, the classical CI optimization criteria suffer from issues such as scale imbalance and correlation mismatch during the fusion process. These deficiencies are not as pronounced in 2D scenarios, where the state space is relatively simple and the observability of each state component is well. However, in 3D scenarios, where the state space is more complex and there are significant disparities in the scale and observability of state components, performance degradation becomes severe. This necessitates the design of specialized mechanisms to improve the data fusion process. In this paper, we identify three main drawbacks of the classical CI optimization criteria in recursive filtering and introduce a weighting mechanism, namely weighted covariance intersection (WCI), to improve its performance. We then introduce WCI into range-based distributed cooperative localization in 3D scenarios, developing a concurrent fusion strategy for multiple distance measurements and designing a weighting matrix based on the error propagation rule of the inertial navigation system (INS). Simulation results demonstrate that the proposed WCI significantly enhances cooperative localization performance compared to classical CI, while the distributed approach outperforms the centralized approach in terms of robustness, scalability, and is more suitable for large-scale swarms.
Index Terms:
Distributed cooperative localization, multi-agent systems, correlation, covariance intersection, weighting mechanism.
I Introduction
Intelligent agents, represented by unmanned aerial vehicles (UAVs) and unmanned ground vehicles (UGVs), have found wide applications in various fields such as environmental monitoring, disaster rescue, and military operations [1, 2, 3]. In recent years, multi-agent systems (MAS) have gained significant attention due to their high flexibility, collaboration capabilities, and robustness, enabling them to efficiently perform complex tasks [4, 5]. The evolution from individual agent intelligence to swarm intelligence has become a key trend in the development of unmanned technologies. Accurate positioning services are the foundation for MAS to execute tasks. Currently, the Global Navigation Satellite System (GNSS) is one of the most widely used positioning methods in MAS [6]. Although GNSS provides all-day and all-weather positioning services in outdoor environments, its signals can be severely obstructed and become unavailable in indoor and underwater scenarios [7]. Moreover, GNSS is not always reliable and is susceptible to interference and spoofing [8]. In contrast, cooperative localization, as an alternative solution, can effectively improve positioning accuracy and availability through relative measurements, information exchange, and data fusion between multiple agents [9].
Since the concept of cooperative localization was first introduced by [10], it has attracted widespread attention and made significant progress over the past three decades. Cooperative localization relies on relative measurements between agents, but these measurements may be correlated with the agents’ states, which must be accounted for during state updates. We refer to the state update process using correlated measurements as correlated update in later illustration. Depending on whether the computation occurs at a central unit or locally at each agent, cooperative localization can be divided into centralized cooperative localization (CCL) and distributed cooperative localization (DCL) [11]. Typically, the two schemes tackle the constantly changing state correlations between agents in different ways. CCL tracks the state correlations between agents with the requirement that all data be collected at a central processing unit, leading to significant communication overhead and poor scalability and robustness. In contrast, DCL typically neither requires nor is capable of tracking state correlations between agents, making it naturally more suitable for large-scale swarms and complex scenarios. However, this also presents the primary challenge of how to perform effective correlated update under unknown state correlations between agents.
Covariance intersection (CI) was initially proposed as a distributed data fusion method capable of producing a consistent and conservative estimate under unknown correlations between data sources [12], and has since been extensively applied to recursive filtering tasks involving unknown correlations between measurements and estimated states, such as target tracking and cooperative perception [13, 14]. Since Arambel first introduced CI to cooperative localization [15], CI and its derivatives have been widely applied in DCL with great success. Carrillo and Li used relative pose observations to obtain two estimates of the agent’s pose and fused them using CI and split covariance intersection (SCI), respectively [11, 16, 17]. Pierre extended CI-based correlated updates to partial observation scenarios, achieving DCL based solely on ranging measurements [18]. Chang also utilized CI to develop a cooperative localization scheme that decouples measurement and communication, improving the robustness and scalability of distributed cooperative localization [19]. However, the above CI-based cooperative localization algorithms are all designed for two-dimensional (2D) scenarios. There is still limited research on DCL for MAS, such as UAV swarms, moving in three-dimensional (3D) environments.
The performance of CI depends on the selection of the weight parameter $\omega$, which is influenced by the optimization criterion. Previous studies have primarily treated CI as a plug-in method for cooperative localization, directly applying classical CI optimization criteria, such as trace or determinant-based metrics of the fused covariance matrix, without considering how different optimization criteria affect the performance of cooperative localization tasks. Furthermore, unlike Kalman filtering (KF), which is inherently a recursive filtering method, CI is a distributed data fusion technique focused on providing consistent estimates in a single-step fusion, and thus exhibits inherent limitations when applied within recursive filtering [20].We investigate the impact of different optimization criteria on cooperative localization performance and identify three major drawbacks of classical CI: scale imbalance, correlation mismatch, and lack of task relevance. Among these, the first two drawbacks are particularly critical, as they can significantly degrade state estimation performance especially when state vector components have large disparities in scale and observability. A detailed discussion of these drawbacks is provided in Section IV. These deficiencies are less pronounced in 2D scenarios, since the state vector consists only of 2D position and orientation, and the scale disparity is relatively small, and the observability of each state component is well. However, in 3D scenarios, where the state space is more complex and there are significant disparities in the scale and observability of state components, performance degradation becomes pronounced. In particular, trace-based optimization in CI performs poorly for small-scale state components, such as attitude and IMU biases, while determinant-based optimization of CI leads to degraded accuracy in position estimation. This highlights the need for specialized mechanisms to improve the correlated update process for DCL in 3D scenarios.
The weighting mechanism is a widely employed method for controlling the data fusion process [21]. Its primary advantage lies in its flexibility and controllability, enabling the adjustment of weights based on specific requirements to achieve optimal fusion results. By incorporating this weighting mechanism into CI, we can improve its performance in correlated data fusion, particularly in balancing state components with significant scale disparities. Furthermore, the weighting mechanism provides a promising strategy for making the optimization criterion task-oriented. In practical applications, different state components can have varying impacts on the final result, and general MSE optimization often fails to accurately reflect the actual contribution of the estimated states to the task at hand. In contrast, weighted mean squared error (WMSE) optimization offers clearer and more relevant significance in these scenarios [22]. Although the concept of weighted covariance intersection (WCI) was first introduced in [15], it primarily focused on enhancing the estimation performance of specific state subspaces and did not highlight the necessity of WCI in scenarios with significant disparities in the scale of state components. Furthermore, it lacked an interpretation of the optimization significance associated with WCI. These two limitations have, to some extent, restricted its application in recursive filtering tasks, such as cooperative localization. In contrast, in this paper, we carefully address these two challenges and successfully integrate WCI into DCL for 3D scenarios, achieving superior performance compared to classical CI.
The main contributions of this paper are summarized as follows:
  • We identify three major drawbacks of CI in recursive filtering: correlation mismatch, scale imbalance, and lack of task relevance. These drawbacks present potential risks in recursive filtering that must be carefully considered and addressed, especially when there are significant differences in the scale and observability of state components. We analyze how WCI can potentially mitigate these drawbacks and clarify its physical meaning in the context of WMSE, as well as its properties.
  • We introduce WCI into range-based distributed cooperative localization in 3D scenarios. We develop a concurrent fusion strategy for multiple distance measurements and design a weighting matrix according to the error propagation rule of INS, addressing the issue of unbalanced and degraded estimation performance between different states in classical CI for this task.
  • We conduct extensive simulation experiments to validate the effectiveness and superiority of the proposed methods. The results show that although WCI does not completely eliminate the inherent drawbacks of CI, it significantly reduces their impact, leading to improved state estimation accuracy in cooperative localization tasks.
The remainder of this paper is organized as follows. In Section II, we introduce the setup of the MAS under study, including the system and state composition, as well as range measurement models. In Section III, we introduce the overall distributed cooperative localization framework and focus our work on improving the correlated update process. Section IV elaborates on the proposed WCI-based correlated update process for DCL in 3D scenarios, including a well-designed weighting matrix selection strategy based on the error propagation rule of INS. In Section V, we present extensive simulations to validate the effectiveness of the proposed DCL method and demonstrate its superiority over CCL method. Finally, the conclusions are summarized in Section VII.
The main notations in the paper are summarized in Table I.
TABLE I: Notation List
Notation Description
lowercase $x$ scalar
bold lowercase $\boldsymbol{x}$ vector
bold uppercase $\boldsymbol{X}$ matrix
$\left\lVert\boldsymbol{x}\right\rVert$ Euclidean norm of a vector
$\mathbb{E}[\cdot]$ expectation operator
$\operatorname{tr}(\cdot)$ trace operator
$\operatorname{det}(\cdot)$ determinant operator
$\operatorname{diag}(\cdot)$ diagonal matrix with the entries inside
$(\cdot)^{\mathsf{T}}$ transpose operator
$\hat{(\cdot)}$ The term with a hat symbol above indicates that it is an estimate contaminated with noise.
$\tilde{(\cdot)}$ The term with a tilde symbol above indicates that it is a measurement contaminated with noise.
$(\cdot)\times$ The antisymmetric matrix constructed from the vector.
$\times$ multiplication cross operator
$\succeq$ Greater than or equal in the Löwner partial order. For example, $\boldsymbol{A}\succeq\boldsymbol{B}$ indicates that $\boldsymbol{A}-\boldsymbol{B}$ is a positive semidefinite matrix.
$\boldsymbol{I}_{M}$ $M\times M$ identity matrix
$\boldsymbol{0}_{M\times N}$ $M\times N$ zero vector or matrix
$\mathbb{R}^{M\times 1}$ $M\times 1$ real-valued vector
II System Setting
We consider a range-based cooperative localization system consisting of a MAS with $N_{u}$ agents moving in 3D space and $N_{a}$ fixed anchors with known positions, as shown in Fig. 1. Since each agent may not be able to connect to a sufficient number of anchors due to communication range limitations or obstructions, it relies on cooperation between agents to achieve accurate state estimation. Each agent is equipped with an onboard IMU to capture its self-motion. In addition, both agents and anchors are equipped with rigidly attached ranging modules. Through signal transceiving, each agent can measure the relative distances to other agents and anchors within its communication range [^1] In this study, anchors are assumed to function solely as signal transmitters, with measurements occurring exclusively on the agent side.. Furthermore, agents and anchors are equipped with communication modules for exchanging information.
Since we use IMU as the proprioceptive sensor, the state of agent $i$ ($i=1,\ldots,N_{a}$) is defined as follows:
Figure 1: Schematic diagram of a range-based cooperative localization system. Due to communication range limitations or obstructions, agents may not be able to obtain sufficient distance measurements with anchors for state estimation. However, through cooperation between agents (including relative distance measurements, information exchange and data fusion), agents can achieve accurate state estimation in such cases.
$\boldsymbol{x}_{i}=\left[\left(\prescript{I}{}{\boldsymbol{p}}_{i}^{n}\right)^{\mathsf{T}},\left(\prescript{I}{}{\boldsymbol{v}}_{i}^{n}\right)^{\mathsf{T}},\left(\boldsymbol{q}_{i}^{n}\right)^{\mathsf{T}},\boldsymbol{b\scriptstyle g}_{i}^{\mathsf{T}},\boldsymbol{b\scriptstyle a}_{i}^{\mathsf{T}}\right]^{\mathsf{T}}$ (1)
where $\prescript{I}{}{\boldsymbol{p}}_{i}^{n}$, $\prescript{I}{}{\boldsymbol{v}}_{i}^{n}$ and $\boldsymbol{q}_{i}^{n}$ represents the position, velocity and orientation (expressed in quaternion form) of the agent in the navigation frame $\mathcal{N}$, respectively. $\boldsymbol{b\scriptstyle g}_{i}$ and $\boldsymbol{b\scriptstyle a}_{i}$ represent the biases of the gyroscope and the accelerometer of the IMU, respectively. The superscript $I$ on the left indicates that the state corresponds to the position of the IMU on the agent. In IMU-based state estimation, system error propagation and state updates are typically performed using the error-state formulation, rather than directly using the nominal state [23]. This approach mitigates the redundancy inherent in quaternion-based attitude representation and addresses the challenges associated with nonlinearity. We define the error-state of agent $i$ as follows:
$\delta\boldsymbol{x}_{i}=\left[\left(\prescript{I}{}{\delta\boldsymbol{p}}_{i}^{n}\right)^{\mathsf{T}},\left(\prescript{I}{}{\delta\boldsymbol{v}}_{i}^{n}\right)^{\mathsf{T}},(\boldsymbol{\phi}^{n}_{i})^{\mathsf{T}},\left(\delta\boldsymbol{b\scriptstyle g}_{i}\right)^{\mathsf{T}},\left(\delta\boldsymbol{b\scriptstyle a}_{i}\right)^{\mathsf{T}}\right]^{\mathsf{T}}$ (2)
Here we adopt Phi-angle model $\boldsymbol{\phi}^{n}_{i}$ to represent the attitude error [24].
In general, the IMU and the ranging module are not co-located on the agent, and there is a lever arm between them. Assuming that each agent has an identical lever arm from the IMU to the ranging module, denoted as $\boldsymbol{l}^{b}$ in the agent’s body-fixed frame $\mathcal{B}$. The relationship between the IMU’s position and the ranging module’s position of agent $i$ can be expressed as follows:
$\prescript{R}{}{\boldsymbol{p}_{i}^{n}}=\prescript{I}{}{\boldsymbol{p}_{i}^{n}}+\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{l}^{b}$ (3)
where $\boldsymbol{C}_{b_{i}}^{n}$ denotes the rotation matrix that transforms a vector from frame $\mathcal{B}_{i}$ to frame $\mathcal{N}$. The relative distances between ranging modules can be expressed as
$d_{ij}^{\mathcal{U}}=\lVert\prescript{R}{}{\boldsymbol{p}_{i}^{n}}-\prescript{R}{}{\boldsymbol{p}_{j}^{n}}\rVert\qquad d_{im}^{\mathcal{A}}=\lVert\prescript{R}{}{\boldsymbol{p}_{i}^{n}}-\prescript{A}{}{\boldsymbol{p}_{m}^{n}}\rVert$ (4)
where $d_{ij}^{\mathcal{U}}$ is the distance between the ranging modules of agent $i$ and agent $j$, while $d_{im}^{\mathcal{A}}$ is the distance between the ranging modules of agent $i$ and anchor $m$ ($\prescript{A}{}{\boldsymbol{p}_{m}^{n}}$ is the position of anchor $m$ expressed in frame $\mathcal{N}$.). The superscripts $\mathcal{U}$ and $\mathcal{A}$ are used to distinguish the relative distances between two agents and those between an agent and an anchor. The distance measurements can be modeled as
${\tilde{d}_{ij}^{\mathcal{U}}}={{d}_{ij}^{\mathcal{U}}}+n_{d_{ij}^{\mathcal{U}}}\qquad{\tilde{d}_{im}^{\mathcal{A}}}={{d}_{im}^{\mathcal{A}}}+n_{d_{im}^{\mathcal{A}}}$ (5)
where $n_{d_{ij}^{\mathcal{U}}}$ and $n_{d_{im}^{\mathcal{A}}}$ are assumed to be i.i.d zero-mean Gaussian with variance $\sigma_{r}^{2}$.
III Distributed Cooperative Localization Framework
In this section, we introduce a range-based distributed cooperative localization framework for 3D scenarios. As illustrated in Fig. 2, we use the state estimation process of agent $i$ at time step $k$ as an example to explain the procedure of the general DCL framework.
In distributed cooperative localization, each agent only estimates its own state and maintains its own state covariance matrix. The state estimation of an agent can be divided into two phases: state prediction and state update. In the state prediction phase, the agent uses its onboard IMU outputs to predict state changes and adjust the covariance matrix for error propagation. In the state update phase, the agent uses distance measurements to other agents and anchors, along with information exchanged with other agents, to correct its state through data fusion methods. More specifically, distance measurements can be categorized into two types: measurements between agents and anchors, and measurements between agents. The former’s noise is independent of agents’ state, while the latter’s is correlated with agents’ state. Therefore, different data fusion methods are required for state updates.
III-A State Prediction Phase
The state prediction phase includes INS mechanism and error propagation, which have been extensively studied and are not the primary focus of our research. Therefore, we will not delve into these aspects here; readers are referred to [25, 23] for further details. It is worth mentioning that during the state prediction phase, states other than position, such as velocity, attitude, and IMU biases, affect the position estimation results. Therefore, for 3D cooperative localization, we must not only focus on the accuracy of position estimation but also ensure a balanced and satisfactory estimation of all states.
Figure 2: Range-based distributed cooperative localization framework for 3D scenarios. Here $\{\boldsymbol{d}_{i}^{\mathcal{U}}\}$ and $\{\boldsymbol{d}_{i}^{\mathcal{A}}\}$ denote the sets of distance measurements from agent $i$ to other agents and anchors, respectively. $\text{info}_{i}$ and $\{\text{info}_{j}\}$ represent the information of agent $i$ that needs to transmitted to other agents, and the information received from other agents, respectively.
III-B State Update Phase
III-B1 Observation Equation
As for the state update phase, we begin with the observation equations for distance measurements. Suppose the predicted state of agent $i$ is defined as follows:
$\hat{\boldsymbol{x}}_{i}=\left[\left(\prescript{I}{}{\hat{\boldsymbol{p}}}_{i}^{n}\right)^{\mathsf{T}},\left(\prescript{I}{}{\hat{\boldsymbol{v}}}_{i}^{n}\right)^{\mathsf{T}},\left(\hat{\boldsymbol{q}}_{i}^{n}\right)^{\mathsf{T}},\hat{\boldsymbol{b\scriptstyle g}}_{i}^{\mathsf{T}},\hat{\boldsymbol{b\scriptstyle a}}_{i}^{\mathsf{T}}\right]^{\mathsf{T}}$
The predicted position of the ranging module can be written as:
$\begin{split}\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}&=\prescript{I}{}{\hat{\boldsymbol{p}}_{i}^{n}}+\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\\ &\approx(\prescript{I}{}{\boldsymbol{p}_{i}^{n}}+{\prescript{I}{}{\delta\boldsymbol{p}_{i}^{n}}})+\left[\boldsymbol{I}-({\boldsymbol{\phi}}_{i}^{n}\times)\right]\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{l}^{b}\\ &=\prescript{R}{}{\boldsymbol{p}_{i}^{n}}+{\prescript{I}{}{\delta\boldsymbol{p}_{i}^{n}}}+[(\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{l}^{b})\times]\boldsymbol{\phi}_{i}^{n}\end{split}$ (6)
Then, the predicted distance between agent $i$ and agent $j$, as well as the predicted distance between agent $i$ and anchor $m$ can be expressed as:
${\hat{d}_{ij}^{\mathcal{U}}}=$ $\left\lVert\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{R}{}{\hat{\boldsymbol{p}}_{j}^{n}}\right\rVert$ (7)
$\approx$ $\left\lVert\prescript{R}{}{\boldsymbol{p}_{i}^{n}}-\prescript{R}{}{\boldsymbol{p}_{j}^{n}}\right\rVert+(\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}})^{\mathsf{T}}\left(\prescript{I}{}{\delta{\boldsymbol{p}}_{i}^{n}}-\prescript{I}{}{\delta{\boldsymbol{p}}_{j}^{n}}\right)$
$+$ $(\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}})^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right]\boldsymbol{\phi}_{i}^{n}-(\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}})^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{j}}^{n}\boldsymbol{l}^{b}\right)\times\right]\boldsymbol{\phi}_{j}^{n}$
${\hat{d}_{im}^{\mathcal{A}}}=$ $\left\lVert\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{A}{}{\boldsymbol{p}}_{m}^{n}\right\rVert$
$\approx$ $\left\lVert\prescript{R}{}{\boldsymbol{p}_{i}^{n}}-\prescript{A}{}{\boldsymbol{p}}_{m}^{n}\right\rVert+(\hat{\boldsymbol{r}}_{im}^{\mathcal{A}})^{\mathsf{T}}\prescript{I}{}{\delta{\boldsymbol{p}}_{i}^{n}}$
$+$ $(\hat{\boldsymbol{r}}_{im}^{\mathcal{A}})^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right]\boldsymbol{\phi}_{i}^{n}$
where $\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}}$ and $\hat{\boldsymbol{r}}_{im}^{\mathcal{A}}$ denote the estimated unit direction vectors pointing from the positions of agent $j$’s and anchor $m$’s ranging modules to the position of agent $i$’s ranging module, respectively. Their explicit expressions are as follows:
$\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}}=\frac{\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{R}{}{\hat{\boldsymbol{p}}_{j}^{n}}}{\left\lVert\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{R}{}{\hat{\boldsymbol{p}}_{j}^{n}}\right\rVert}\qquad\hat{\boldsymbol{r}}_{im}^{\mathcal{A}}=\frac{\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{A}{}{{\boldsymbol{p}}_{m}^{n}}}{\left\lVert\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{A}{}{{\boldsymbol{p}}_{m}^{n}}\right\rVert}$ (8)
Since we use the error-state formulation for error propagation and state updates, we apply differential measurements rather than the original measurements during the state update phase. The differential measurements, i.e., the differences between the predicted distances and actual distance measurements, can be expressed as:
$\delta{d_{ij}^{\mathcal{U}}}$ $={\hat{d}_{ij}^{\mathcal{U}}}-{\tilde{d}_{ij}^{\mathcal{U}}}=\boldsymbol{H}_{ij}^{\mathcal{U}}\delta\boldsymbol{x}_{i}+\boldsymbol{G}_{ij}^{\mathcal{U}}\boldsymbol{v}_{ij}^{\mathcal{U}}$ (9)
$\delta{d_{im}^{\mathcal{A}}}$ $={\hat{d}_{im}^{\mathcal{A}}}-{\tilde{d}_{im}^{\mathcal{A}}}=\boldsymbol{H}_{ij}^{\mathcal{A}}\delta\boldsymbol{x}_{i}+v_{im}^{\mathcal{A}}$
Here, $\boldsymbol{v}_{ij}^{\mathcal{U}}$ and $v_{im}^{\mathcal{A}}$ are the noise of the differential measurements, where $v_{im}^{\mathcal{A}}=-n_{d_{im}^{\mathcal{A}}}$, and the expressions of $\boldsymbol{v}_{ij}^{\mathcal{U}}$ and its corresponding covariance matrix $\boldsymbol{R}_{v_{ij}}$ are given by
$\boldsymbol{v}_{ij}^{\mathcal{U}}=\left[\left(\prescript{I}{}{\delta{\boldsymbol{p}}_{j}^{n}}\right)^{\mathsf{T}},\left(\boldsymbol{\phi}_{j}^{n}\right)^{\mathsf{T}},n_{d_{ij}^{\mathcal{U}}}\right]^{\mathsf{T}}$ (10)
$\boldsymbol{R}_{ij}^{\mathcal{U}}=\begin{bmatrix}&\boldsymbol{P}_{{p}_{j}},&\boldsymbol{P}_{{p}_{j}{\phi}_{j}},&\boldsymbol{0}_{3\times 1}\\ &\boldsymbol{P}_{{p}_{j}{\phi}_{j}}^{\mathsf{T}},&\boldsymbol{P}_{{\phi}_{j}},&\boldsymbol{0}_{3\times 1}\\ &\boldsymbol{0}_{1\times 3},&\boldsymbol{0}_{1\times 3},&\sigma_{r}^{2}\end{bmatrix}$ (11)
Here $\boldsymbol{P}_{{p}_{j}}$, $\boldsymbol{P}_{{\phi}_{j}}$, $\boldsymbol{P}_{{p}_{j}{\phi}_{j}}$ are the submatrices of the state covariance matrix of agent $j$, corresponding to the covariance matrix of its position, attitude, and their cross-covariance, respectively. Additionally, the expressions of $\boldsymbol{H}_{ij}^{\mathcal{U}}$, $\boldsymbol{G}_{ij}^{\mathcal{U}}$ and $\boldsymbol{H}_{ij}^{\mathcal{A}}$ are given by
$\boldsymbol{H}_{ij}^{\mathcal{U}}$ $=\left[\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}}\right)^{\mathsf{T}},\boldsymbol{0}_{3\times 3},\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}}\right)^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right],\boldsymbol{0}_{3\times 6}\right]$ (12)
$\boldsymbol{G}_{ij}^{\mathcal{U}}$ $=\left[-\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}}\right)^{\mathsf{T}},-\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{U}}\right)^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right],-1\right]$
$\boldsymbol{H}_{im}^{\mathcal{A}}$ $=\left[\left(\hat{\boldsymbol{r}}_{im}^{\mathcal{A}}\right)^{\mathsf{T}},\boldsymbol{0}_{3\times 3},\left(\hat{\boldsymbol{r}}_{im}^{\mathcal{A}}\right)^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right],\boldsymbol{0}_{3\times 6}\right]$
Assume that agent $i$ obtains the distances to $M$ anchors and $N$ other agents for state updates at time step $k$. Without loss of generality, let the indices of the $M$ anchors be $1$ to $M$, and the indices of the $N$ other agents be $1$ to $N$. Let the measurement vector formed by the differential distances to the $M$ anchors be denoted as $\boldsymbol{z}_{i}^{\mathcal{A}}=\left[\delta d_{i1}^{\mathcal{A}},\cdots,\delta d_{iM}^{\mathcal{A}}\right]^{\mathsf{T}}\in\mathbb{R}^{M\times 1}$, with the corresponding covariance matrix being $\boldsymbol{R}_{i}^{\mathcal{A}}=\sigma_{r}^{2}\boldsymbol{I}_{M}$; Let the differential distances to the $N$ other agents and their corresponding covariance matrices be denoted as $\left\{\delta d_{ij}^{\mathcal{U}},\boldsymbol{R}_{ij}^{\mathcal{U}}\right\}$, $j=1,\ldots,N$, respectively. The predicted state and the error-state covariance matrix of agent $i$ before the update are given by $\left\{\hat{\boldsymbol{x}}_{i,k|k-1},\boldsymbol{P}_{i,k|k-1}\right\}$.
III-B2 Independent Update
The distance measurements between agent $i$ and anchors, which are independent of the agent’s state, can be directly used for state update via EKF. Since the distances to the $M$ anchors are mutually independent, there is no difference between updating them sequentially as individual observations or combining them into a single observation vector for concurrent updating. We choose the latter approach, where all anchor measurements are aggregated into a single observation vector for the update. The corresponding update procedure can be written as:
$\boldsymbol{H}_{i}^{\mathcal{A}}=[{\boldsymbol{H}_{i1}^{\mathcal{A}}}^{\mathsf{T}},\cdots,{\boldsymbol{H}_{iM}^{\mathcal{A}}}^{\mathsf{T}}]^{\mathsf{T}}$
$\boldsymbol{K}=\boldsymbol{P}_{i,k|k-1}{\boldsymbol{H}_{i}^{\mathcal{A}}}^{\mathsf{T}}\left(\boldsymbol{H}_{i}^{\mathcal{A}}\boldsymbol{P}_{i,k|k-1}{\boldsymbol{H}_{i}^{\mathcal{A}}}^{\mathsf{T}}+\boldsymbol{R}_{i}^{\mathcal{A}}\right)^{-1}$
$\delta\boldsymbol{x}_{i,k}=\boldsymbol{K}\boldsymbol{z}_{i}^{\mathcal{A}}$ (13)
$\boldsymbol{P}_{i,k}=\left(\boldsymbol{I}-\boldsymbol{K}\boldsymbol{H}_{i}^{\mathcal{A}}\right)\boldsymbol{P}_{i,k|k-1}$
The estimated error-state $\delta\boldsymbol{x}_{i,k}$ can then be used to correct the estimated nominal state $\hat{\boldsymbol{x}}_{i,k|k-1}$, resulting in the updated nominal state $\hat{\boldsymbol{x}}_{i,k}$.
III-B3 Correlated Update
For distance measurements between agents, consider the measurement between agent $i$ and agent $j$ as an example, as shown in (9) and (10). The differential distance $\delta d_{ij}^{\mathcal{U}}$ is used for updating agent $i$’s state, and its noise includes the state estimation error of agent $j$. This leads to potential correlations between the differential distance measurement and agent $i$’s state. In a distributed framework, tracking the state correlation between agent $i$ and agent $j$ is extremely challenging or even infeasible. If the correlation is ignored and assumed to be independent, it could lead to the “data incest” problem [26], resulting in state estimation divergence. Therefore, our work focuses on this correlated update process, with the goal of achieving effective state updates using relative distance measurements between agents under unknown correlations, ensuring accurate and consistent state estimation in 3D cooperative localization scenarios.
IV Methodology
In this section, we present our methods for DCL in 3D scenarios. We begin by revisiting CI and discussing three primary drawbacks of the classical CI criteria in recursive filtering. Next, we introduce a weighting mechanism, namely WCI, to mitigate the negative impacts of CI and analyze its properties. Finally, we propose a WCI-based approach for correlated updates in DCL, which accounts for the characteristics of CI fusion and the error propagation rule of INS, balancing the estimation of all state components with scale disparity.
IV-A Covariance Intersection Revisit
IV-A1 Formulas of Covariance Intersection
Consider a state vector $\boldsymbol{x}$ with two unbiased and consistent estimates $\{\hat{\boldsymbol{{x}}}_{1},\boldsymbol{P}_{1}\}$ and $\{\hat{\boldsymbol{{x}}}_{2},\boldsymbol{P}_{2}\}$. Each estimate can be written as $\hat{\boldsymbol{{x}}}_{i}=\boldsymbol{x}+\delta{\boldsymbol{x}}_{i}$, $i=1,2$, where $\mathbb{E}[\delta{\boldsymbol{x}}_{i}]=\boldsymbol{0}$. We define the actual covariance matrices and the cross covariance matrix as
$\boldsymbol{P}_{i}^{*}=\mathbb{E}\left[\delta{\boldsymbol{x}}_{i}\delta{\boldsymbol{x}}_{i}^{\mathsf{T}}\right]\quad\boldsymbol{P}_{12}^{*}=\mathbb{E}\left[\delta{\boldsymbol{x}}_{1}\delta{\boldsymbol{x}}_{2}^{\mathsf{T}}\right]$ (14)
Since $\boldsymbol{P}_{i}^{*}$ are generally unknown, we name $\boldsymbol{P}_{1}$ and $\boldsymbol{P}_{2}$ as nominal covariance matrices. Consistency of each estimate then requires $\boldsymbol{P}_{i}\succeq\boldsymbol{P}_{i}^{*}$ [27].
CI produces an unbiased and consistent fused estimate $\{\hat{\boldsymbol{{x}}},\boldsymbol{P}\}$ without the knowledge of cross-correlation information as [12]
$\boldsymbol{P}$ $=\left(\omega\boldsymbol{P}_{1}^{-1}+(1-\omega)\boldsymbol{P}_{2}^{-1}\right)^{-1}$ (15)
$\hat{\boldsymbol{{x}}}$ $=\boldsymbol{\boldsymbol{P}}\left(\omega\boldsymbol{P}_{1}^{-1}\hat{\boldsymbol{x}}_{1}+(1-\omega)\boldsymbol{P}_{2}^{-1}\hat{\boldsymbol{x}}_{2}\right)$
where $\omega\in[0,1]$ is called the weight parameter. It has been proved that for any admissible $\omega$ and $\boldsymbol{P}_{12}^{*}$, the fused covariance matrix satisfies $\boldsymbol{P}\succeq\boldsymbol{P}^{*}$. The weight parameter $\omega$ is selected by minimizing a certain cost function $f(\boldsymbol{P};\omega)$:
$\omega^{*}=\arg\min_{\omega\in[0,1]}f(\boldsymbol{P};\omega)$ (16)
Typically, the cost function is chosen as the trace or determinant of the fused covariance matrix, i.e., $f(\boldsymbol{P})=\operatorname{tr}(\boldsymbol{P})$ or $\text{det}(\boldsymbol{P})$; we refer to them as the classical criteria.
Furthermore, for partial-observation scenarios like $\boldsymbol{x}_{2}=\boldsymbol{H}\boldsymbol{x}$, where $\boldsymbol{H}$ is rank-deficient, and $\boldsymbol{P}_{1}^{-1}+\boldsymbol{H}^{\mathsf{T}}\boldsymbol{P}_{2}^{-1}\boldsymbol{H}$ is invertible, the CI estimate is given by [15]
$\boldsymbol{P}=\left(\omega\boldsymbol{P}_{1}^{-1}+(1-\omega)\boldsymbol{H}^{\mathsf{T}}\boldsymbol{P}_{2}^{-1}\boldsymbol{H}\right)^{-1}$ (17)
$\hat{\boldsymbol{{x}}}=\boldsymbol{\boldsymbol{P}}\left(\omega\boldsymbol{P}_{1}^{-1}\hat{\boldsymbol{x}}_{1}+(1-\omega)\boldsymbol{H}^{\mathsf{T}}\boldsymbol{P}_{2}^{-1}\hat{\boldsymbol{x}}_{2}\right)$
IV-A2 Drawbacks of the Classical Criteria
Although the classical CI criteria have achieved widespread success in various applications, they can degrade performance in certain scenarios. We identify three primary drawbacks of classical criteria in the context of recursive filtering:
  • Scale imbalance
    When state components differ substantially in units or magnitudes, the trace-based criterion becomes dominated by the larger-scale components and tends to overly magnify the uncertainty of small-scale components. We give a simple two-dimensional example to demonstrate it.
    Let $\boldsymbol{P}_{1}=\begin{bmatrix}1&0\\ 0&0.01\end{bmatrix}$ and $\boldsymbol{P}_{2}=\begin{bmatrix}1.09&0\\ 0&0.001\end{bmatrix}$. Although $\boldsymbol{P}_{2}$ reduces the second component’s variance by 90% dramatically, its 9% increase in the first component drives the trace-minimization fusion (CI-trace) to favor $\boldsymbol{P}_{1}$. As shown in Fig. 3(a), the fused covariance matrix thus remains almost identical to $\boldsymbol{P}_{1}$, offering negligible improvement for the small-scale component. By contrast, determinant-based criterion (CI-det) favors $\boldsymbol{P}_{2}$ because the determinant corresponds to the uncertainty volume in which components with small scales can play a role. However, both criteria lack the ability to flexibly adjust and balance the estimation performance of different components in the presence of significant scale disparity.
  • Correlation mismatch
    The actual covariance matrix of the CI estimate, given by (18), depends on the actual cross‐covariance $\boldsymbol{P}_{12}^{*}$, whereas the nominal covariance matrix $\boldsymbol{P}$ in (15) is completely independent of cross‐correlations. This ensures consistency by guaranteeing that the nominal covariance matrix dominates the actual covariance matrix in Löwner partial order for all admissible cross-correlations, but offers no alignment of intra-state correlations. When the actual correlations diverge substantially from those implied by $\boldsymbol{P}$, recursive filtering can experience severe error divergence, particularly under partial observation or scale disparity cases.
    $\begin{split}\boldsymbol{P}^{*}&=\mathbb{E}\left\{\delta{\boldsymbol{x}}\delta{\boldsymbol{x}}^{\mathsf{T}}\right\}\\ &=\boldsymbol{P}[\ \omega^{2}\boldsymbol{P}_{1}^{-1}\boldsymbol{P}_{1}^{*}\boldsymbol{P}_{1}^{-1}+(1-\omega)\omega\boldsymbol{P}_{1}^{-1}\boldsymbol{P}_{12}^{*}\boldsymbol{P}_{2}^{-1}+\\ &(1-\omega)\omega\boldsymbol{P}_{2}^{-1}\boldsymbol{P}_{12}^{*\mathsf{T}}\boldsymbol{P}_{1}^{-1}+(1-\omega)^{2}\boldsymbol{P}_{2}^{-1}\boldsymbol{P}_{2}^{*}\boldsymbol{P}_{2}^{-1}\ ]\boldsymbol{P}\end{split}$ (18)
    For example, consider two estimates of $\boldsymbol{x}$: the first estimate is full observed with $\boldsymbol{P}_{1}=\begin{bmatrix}1&0.008\\ 0.008&0.0001\end{bmatrix}$, while the second estimate is partially observed with $\boldsymbol{H}=[1,0]$ and $\boldsymbol{P}_{2}=0.25$. Without loss of generality, we assume that the two estimates are uncorrelated—though this is unknown a priori—so we apply CI to fuse them. Fig. 3(b) shows that the CI‐fused nominal covariance ellipse encloses the actual covariance ellipse, confirming CI’s consistency and conservativeness. However, the nominal correlation coefficient given by CI-trace, $\rho_{1}=0.08$, deviates drastically from the actual correlation coefficient, $\rho_{1}^{*}=0.55$, which can lead to unsatisfactory updates in subsequent recursive filtering. In contrast, the correlation deviation of CI-det is relatively small but still exists in this example, suggesting that it is feasible to select an appropriate optimization criterion to mitigate the correlation mismatch problem.
(a) Example for scale imbalance
(b) Example for correlation mismatch
Figure 3: Covariance ellipses before and after CI fusion (axes scaled differently for better demonstration). (a) The covariance ellipse of CI-trace almost overlaps that of $\boldsymbol{P}_{1}$, showing that fusion is dominated by the larger-scale state component—so the fused estimate closely matches the lower-uncertainty estimate along that dimension. While CI-det can account for smaller-scale state component, but still lacks balance between two components. (b) The nominal covariance ellipse encloses the actual covariance ellipse but differs markedly in shape and orientation, reflecting a significant mismatch between implied and actual inter-component correlations.
  • Lack of task relevance
    As Hurley pointed out, the classical determinant-based criterion corresponds to minimizing the Shannon information of the fused density function [28]. However, the perspective of information theory offers little interpretation of performance in physical world. Meanwhile, the classical trace-based criterion implements a minimum MSE objective. However, when the state vector comprises heterogeneous components (e.g., position versus orientation), the aggregated MSE lacks a clear physical interpretation as well. In summary, both the two classical criteria ignores the task‐dependent significance of individual components, rendering them too coarse for applications that require prioritization among state variables.
IV-B Weighted Covariance Intersection
The classical trace-based criterion is derived from the principle of minimizing MSE of the estimate—equivalently, the trace of the actual covariance matrix (cf. (19)). In practice, since the actual covariance matrix is unavailable, the trace of the nominal covariance matrix is minimized instead.
$\min\mathbb{E}[(\boldsymbol{x}-\boldsymbol{\hat{x}})^{\mathsf{T}}(\boldsymbol{x}-\boldsymbol{\hat{x}}))]=\operatorname{tr}(\boldsymbol{P}^{*})\rightarrow\min\operatorname{tr}(\boldsymbol{P})$ (19)
As noted above, minimizing MSE alone is too coarse and lacks task specificity. To address this, Arambel et al. introduced relative weights to control the estimation process [15], termed WCI. However, they treated these weights merely as tuning knobs to improve the accuracy of a chosen component, without providing a theoretical rationale. We show in (20) that minimizing the weighted trace of the fused covariance matrix is mathematically equivalent to minimizing a WMSE over the entire state. By selecting the weighting matrix $\boldsymbol{W}$ to reflect task‐dependent priorities or target WMSE levels, WCI reintroduces task relevance to the fusion criterion.
$\min\mathbb{E}[(\boldsymbol{x}-\boldsymbol{\hat{x}})^{\mathsf{T}}\boldsymbol{W}(\boldsymbol{x}-\boldsymbol{\hat{x}}))]=\operatorname{tr}(\boldsymbol{W}\boldsymbol{P}^{*})\rightarrow\min\operatorname{tr}(\boldsymbol{W}\boldsymbol{P})$ (20)
Moreover, by assigning greater weight to smaller-scale components, the weighting mechanism helps mitigate scale imbalance and ensures a more balanced fusion across all state variables.
Last but not least, WCI provides additional flexibility in selecting the fusion weight $\omega$, which may help mitigate correlation mismatch. Although the actual intra‐state correlations are unpredictable due to the unknown cross-correlation in (18), in the limiting cases $\omega=0$ or $1$ the cross‐correlation term vanishes from the fusion equations, ensuring that the unknown cross-correlations do not affect the final fused estimate. Thus, selecting $\omega$ closer to 0 or $1$ may reduce the potential deviation between the actual covariance matrix and its nominal counterpart.
In addition, we identify two properties of WCI: consistency and optimality, which are summarized in Proposition 1 and Proposition 2, respectively.
Proposition 1.
(Consistency) For any weight matrix, WCI provides consistent estimates.
Proof.
Because CI provides consistent estimates for any $\omega\in[0,1]$, and WCI only changes the selection of $\omega$ which is still within $[0,1]$, the resulting estimates remain consistent. ∎
Proposition 2.
(Optimality) For any positive-definite weight matrix, WCI provides the optimal bound for the fusion of two estimates under unknown correlations in terms of WMSE.
The proof of Proposition 2 is provided in Appendix.
In summary, WCI effectively mitigates the three primary drawbacks of the classical criteria of CI in recursive filtering. By selecting task‐specific weighting matrices, practitioners can tailor CI to their needs, potentially achieving improved estimation performance.
IV-C WCI-based Correlated Update in DCL
In this subsection, continuing from the distributed cooperative localization framework introduced in Section III, we present the WCI-based correlated update process. Using agent $i$ as an example, assume it has obtained differential distances to the $N$ other agents and their corresponding covariance matrices be denoted as $\left\{\delta d_{ij}^{\mathcal{U}},\boldsymbol{R}_{ij}^{\mathcal{U}}\right\}$, $j=1,\ldots,N$, whose explicit expressions are given in (9) and (11), respectively. Since the noise of the differential distance measurements contains estimation errors of other agents’ states, which are correlated with the state of agent $i$, and these correlations are unknown, we can use CI, or its enhanced version WCI, introduced in Section III-B, to perform the correlated update process.
Unlike EKF, where sequential and concurrent updates of multiple observations yield the same result, CI does not provide the same guarantee. In fact, CI ensures that the fusion of two data sources provides an optimal conservative upper bound, but it cannot guarantee the same for the fusion of multiple data sources [29]. Sequential updates of multiple observations lead to additional performance degradation and result in a more conservative upper bound. Therefore, for multiple distance measurements between agents, we combine them into a single observation vector for state updates.
$\boldsymbol{z}_{i}^{\mathcal{U}}=\left[\delta{d_{i1}^{\mathcal{U}}},\cdots,\delta{d_{iN}^{\mathcal{U}}}\right]^{\mathsf{T}}=\boldsymbol{H}_{i}^{\mathcal{U}}\delta\boldsymbol{x}^{i}+\boldsymbol{G}_{i}^{\mathcal{U}}\boldsymbol{v}_{i}^{\mathcal{U}}$ (21)
where the explicit expressions of $\boldsymbol{H}_{i}^{\mathcal{U}}$ and $\boldsymbol{G}_{i}^{\mathcal{U}}$ are given by:
$\boldsymbol{H}_{i}^{\mathcal{U}}$ $=[{\boldsymbol{H}_{i1}^{\mathcal{U}}}^{\mathsf{T}},\cdots,{\boldsymbol{H}_{iN}^{\mathcal{U}}}^{\mathsf{T}}]^{\mathsf{T}}$ (22)
$\boldsymbol{G}_{i}^{\mathcal{U}}$ $=[{\boldsymbol{G}_{i1}^{\mathcal{U}}}^{\mathsf{T}},\cdots,{\boldsymbol{G}_{iN}^{\mathcal{U}}}^{\mathsf{T}}]^{\mathsf{T}}$
The combined noise vector and its actual covariance matrix can be written as:
$\boldsymbol{v}_{i}^{\mathcal{U}}$ $=\left[{\boldsymbol{v}_{i1}^{\mathcal{U}}}^{\mathsf{T}},\cdots,{\boldsymbol{v}_{iN}^{\mathcal{U}}}^{\mathsf{T}}\right]^{\mathsf{T}}$ (23)
$\boldsymbol{R}_{i}^{\mathcal{U}}$ $=\begin{bmatrix}\boldsymbol{R}_{i1}^{\mathcal{U}}&\cdots&\boldsymbol{C}_{i,1N}^{\mathcal{U}}\\ \vdots&\ddots&\vdots\\ {\boldsymbol{C}_{i,1N}^{\mathcal{U}}}^{\mathsf{T}}&\cdots&\boldsymbol{R}_{iN}^{\mathcal{U}}\end{bmatrix}$
Here, $\boldsymbol{C}_{i,1N}^{\mathcal{U}}$ denote the cross-covariance matrix between $\boldsymbol{v}_{i1}^{\mathcal{U}}$ and $\boldsymbol{v}_{iN}^{\mathcal{U}}$, which is generally non-zero since the states of agent 1 and agent $N$ are potentially correlated. However, $\boldsymbol{C}_{i,1N}^{\mathcal{U}}$ is not unavailable in practical distributed systems, as we do not track the correlations between agents. An alternative approach is to appropriately inflate the noise covariance matrix while ensuring that the nominal covariance matrix remains greater than the actual covariance matrix, thus maintaining consistency. We focus on the diagonal elements of $\boldsymbol{R}_{i}^{\mathcal{U}}$ and multiply the agent state error variance terms by 2 to obtain the nominal noise covariance matrix $\boldsymbol{R}_{i}^{\mathcal{U}^{\prime}}$, as follows:
$\boldsymbol{R}_{i}^{\mathcal{U}^{\prime}}=\begin{bmatrix}2\boldsymbol{P}_{p_{1}}^{d}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}&\cdots&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{3\times 3}&2\boldsymbol{P}_{\phi_{1}}^{d}&\boldsymbol{0}_{3\times 1}&\cdots&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&\sigma_{r}^{2}&\cdots&\boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}&\cdots&2\boldsymbol{P}_{p_{N}}^{d}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}&\cdots&\boldsymbol{0}_{3\times 3}&2\boldsymbol{P}_{\phi_{N}}^{d}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&0&\cdots&\boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&\sigma_{r}^{2}\end{bmatrix}$ (24)
where
$\boldsymbol{P}_{p_{j}}^{d}=\operatorname{diag}(\left[\sigma_{p_{j,x}}^{2},\sigma_{p_{j,y}}^{2},\sigma_{p_{j,z}}^{2}\right]^{\mathsf{T}})$ (25)
$\boldsymbol{P}_{\phi_{j}}^{d}=\operatorname{diag}(\left[\sigma_{\phi_{j,x}}^{2},\sigma_{\phi_{j,y}}^{2},\sigma_{\phi_{j,z}}^{2}\right]^{\mathsf{T}})$
($j=1,\ldots,N$) are the diagonal matrices consisting of the error variances of agent $j$’s position and attitude. It is easy to observe that $\boldsymbol{R}_{i}^{\mathcal{U}^{\prime}}\succeq\boldsymbol{R}_{i}^{\mathcal{U}}$, meaning that the nominal covariance matrix $\boldsymbol{R}_{i}^{\mathcal{U}^{\prime}}$ still provides a consistent estimate of the differential measurement vector. In addition, the nominal covariance matrix is free from the correlation items, thereby eliminating the need to track correlations and reducing the communication burden.
With the differential measurements between agents $\left\{\boldsymbol{z}_{i}^{\mathcal{U}},\boldsymbol{R}_{i}^{\mathcal{U}^{\prime}}\right\}$, we can apply CI for state updates. A KF-style CI formulation for partial observations is provided in [16] as follows:
$\boldsymbol{P}_{1}=\frac{1}{\omega}\boldsymbol{P}_{i,k|k-1}$
$\boldsymbol{P}_{2}=\frac{1}{1-\omega}\boldsymbol{R}_{i}^{\mathcal{U}^{\prime}}$
$\boldsymbol{K}=\boldsymbol{P}_{1}{\boldsymbol{H}_{i}^{\mathcal{U}}}^{\mathsf{T}}\left({\boldsymbol{H}_{i}^{\mathcal{U}}}\boldsymbol{P}_{1}{\boldsymbol{H}_{i}^{\mathcal{U}}}^{\mathsf{T}}+\boldsymbol{P}_{2}\right)^{-1}$ (26)
$\delta{\boldsymbol{x}}_{i,k}=\boldsymbol{K}\boldsymbol{z}_{i}^{\mathcal{U}}$
$\boldsymbol{P}_{i,k}=(\boldsymbol{I}-\boldsymbol{K}{\boldsymbol{H}_{i}^{\mathcal{U}}})\boldsymbol{P}_{1}$
Here, $\omega$ is the weight parameter, which is determined by a specific optimization criterion in (16). We find that both the classical criteria of minimizing the trace or the determinant of the fused covariance matrix perform poorly in this problem. Instead, we choose WCI as presented in (20), with the primary challenge lying in the selection of the weighting matrix.
$\boldsymbol{S}=\begin{bmatrix}\boldsymbol{I}_{3},\ \boldsymbol{I}_{3}\cdot T_{a},\ \frac{1}{2}\left[\left(\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{f}_{i}^{b}\right)\times\right]T_{a}^{2},\ -\frac{1}{6}\left[\left(\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{f}_{i}^{b}\right)\times\right]\boldsymbol{C}_{b_{i}}^{n}T_{a}^{3},\ \frac{1}{2}\boldsymbol{C}_{b_{i}}^{n}T_{a}^{2}\end{bmatrix}$ (30)
$\mathbb{E}\left[\left(\prescript{I}{}{\delta{\boldsymbol{p}}}_{i,T_{a}}^{n}\right)^{\mathsf{T}}\left(\prescript{I}{}{\delta{\boldsymbol{p}}}_{i,T_{a}}^{n}\right)\right]=\operatorname{tr}\left(\mathbb{E}\left[\left(\prescript{I}{}{\delta{\boldsymbol{p}}}_{i,T_{a}}^{n}\right)\left(\prescript{I}{}{\delta{\boldsymbol{p}}}_{i,T_{a}}^{n}\right)^{\mathsf{T}}\right]\right)=\operatorname{tr}\left(\boldsymbol{S}\mathbb{E}\left[\delta\boldsymbol{x}_{i,0}\delta\boldsymbol{x}_{i,0}^{\mathsf{T}}\right]\boldsymbol{S}^{\mathsf{T}}\right)=\operatorname{tr}\left(\boldsymbol{S}^{\mathsf{T}}\boldsymbol{S}\boldsymbol{P}_{i,0}\right)$ (31)
In the complete state estimation procedure for agents, inertial navigation serves as the primary method for state propagation, while distance measurements are used as a supplementary tool for state updates. If the error in inertial navigation can be minimized as much as possible, the overall state estimation error in cooperative localization will decrease, thereby enhancing the agent’s ability to navigate autonomously in scenarios with insufficient distance measurements. In summary, the optimization goal in the state update phase can be framed as minimizing the inertial navigation error.
We consider the propagation law of inertial navigation errors. Generally, agents are equipped with low cost micro-electromechanical system (MEMS) sensors, so we do not account for factors such as the earth’s rotation, the variation of frame $\mathcal{N}$, or changes in gravity. As a result, a simplified error state model can be expressed as [24]:
$\left\{\begin{aligned} \prescript{I}{}{\delta{\dot{\boldsymbol{p}}}_{i}^{n}}&=\prescript{I}{}{\delta{\boldsymbol{v}}_{i}^{n}}\\ \prescript{I}{}{\delta{\dot{\boldsymbol{v}}}_{i}^{n}}&=\boldsymbol{f}_{i}^{n}\times\boldsymbol{\phi}_{i}^{n}+\boldsymbol{C}_{b_{i}}^{n}\delta\boldsymbol{f}_{i}^{b}\\ \dot{\boldsymbol{\phi}}_{i}^{n}&=-\boldsymbol{C}_{b_{i}}^{n}\delta\boldsymbol{\omega}_{i}^{b}\end{aligned}\right$. (27)
where $\boldsymbol{f}_{i}^{n}=\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{f}_{i}^{b}$ is the projection of the specific force $\boldsymbol{f}_{i}^{b}$ into frame $\mathcal{N}$. $\delta\boldsymbol{f}_{i}^{b}$ and $\delta\boldsymbol{\omega}_{i}^{b}$ represent the biases in the IMU outputs, specifically the specific force and angular velocity, with respect to their actual values. These biases are modeled as:
$\delta\boldsymbol{f}_{i}^{b}=\boldsymbol{b\scriptstyle a}_{i}+\boldsymbol{w}_{a}$ (28)
$\delta\boldsymbol{\omega}_{i}^{b}=\boldsymbol{b\scriptstyle g}_{i}+\boldsymbol{w}_{g}$
where $\boldsymbol{w}_{a}$ and $\boldsymbol{w}_{g}$ are zero-mean random noise. The error state model indicates that the accumulation of inertial navigation position errors is related to the initial position error by a zero-order relationship, which remains constant over time. It is related to the initial velocity error by a first-order relationship, which increases linearly with time. The position error is related to the initial attitude and acceleration errors by a second-order relationship, which increases quadratically with time, and it is related to the initial gyroscope bias error by a third-order relationship, which increases cubically with time[25].
We make a simple assumption that inertial navigation position error accumulates similarly to the case where agent $i$ moves with constant velocity in a straight line over a time period of $T_{a}$. The inertial navigation position error at time $T_{a}$ can then be approximately written as:
$\begin{split}\prescript{I}{}{\delta{\boldsymbol{p}}}_{i,T_{a}}^{n}=\ &\prescript{I}{}{\delta{\boldsymbol{p}}}_{i,0}^{n}+\prescript{I}{}{\delta{\boldsymbol{v}}}_{i,0}^{n}T_{a}+\frac{1}{2}\left[\left(\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{f}^{b}\right)\times\right]\boldsymbol{\phi}_{i}^{n}T_{a}^{2}+\\ &\frac{1}{2}\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{b\scriptstyle a}_{i}T_{a}^{2}-\frac{1}{6}\left[\left(\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{f}^{b}\right)\times\right]\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{b\scriptstyle g}_{i}T_{a}^{3}\\ =\ &\boldsymbol{S}\delta\boldsymbol{x}_{i,0}\end{split}$ (29)
where the explicit expression for $\boldsymbol{S}$ is given by (30). We aim to minimize the mean square error of $\prescript{I}{}{\delta{\boldsymbol{p}}}_{i,T_{a}}^{n}$, which corresponds to minimizing the trace of the product of $\boldsymbol{S}^{\mathsf{T}}\boldsymbol{S}$ and initial error-state covariance matrix $\boldsymbol{P}_{i,0}$ as illustrated in (31).
It’s interesting to notice that if we assign the weighting matrix $\boldsymbol{W}$ as
$\boldsymbol{W}=\boldsymbol{S}^{\mathsf{T}}\boldsymbol{S}$ (32)
The optimization goal in (31) is the same as that in (20). We can finally determine the weight parameter $\omega$ in (26) by
$\omega^{*}=\arg\min_{\omega\in[0,1]}\operatorname{tr}(\boldsymbol{W}_{i,k}\boldsymbol{P}_{i,k})$ (33)
where $\boldsymbol{W}_{i,k}$ is determined by (30) and (32) with predicted state estimates and IMU outputs at time step $k$. The estimated error-state $\delta\boldsymbol{x}_{i,k}$ can be used to correct nominal state estimates as well. The entire procedures of the WCI-based correlated update are summarized in Algorithm 1.
Algorithm 1 WCI-based Correlated Update
1:
2:
  • State estimate after prediction $\{\hat{\boldsymbol{x}}_{i,k|k-1},\boldsymbol{P}_{i,k|k-1}\}$
  • IMU measurement $\tilde{\boldsymbol{f}}_{k,i}^{b}$
  • Distance measurements with other agents $\{d_{ij}^{\mathcal{U}}\}$
  • Other agents’ position and attitude estimates $\prescript{\mathcal{I}}{}{\hat{\boldsymbol{p}}_{j,k|k-1}^{n}}$ and $\hat{\boldsymbol{q}}_{j,k|k-1}^{n}$, along with their nominal variances $\{\sigma_{p_{j,x}}^{2},\sigma_{p_{j,y}}^{2}\sigma_{p_{j,z}}^{2}\}$ and $\{\sigma_{\phi_{j,x}}^{2},\sigma_{\phi_{j,y}}^{2}\sigma_{\phi_{j,z}}^{2}\}$.
3:State estimate after update $\{\hat{\boldsymbol{x}}_{i,k},\boldsymbol{P}_{i,k}\}$
4:Calculate differential measurements with the inflated nominal covariance matrices $\{\boldsymbol{z}_{i}^{\mathcal{U}},\boldsymbol{R}_{i}^{\mathcal{U}^{\prime}}\}$ according to (17), (21), (24) and (25).
5:Determine the weighting matrix $\boldsymbol{W}_{i,k}$ by (30) and (32), replacing ${\boldsymbol{f}}_{k,i}^{b}$ with $\tilde{\boldsymbol{f}}_{k,i}^{b}$.
6:Execute data fusion using WCI by (26) and (33).
7:Perform error feedback to correct the predicted state estimate.
Furthermore, we have described the independent and correlated update procedures for distance measurements between agents and anchors, and between agents, using EKF and WCI, respectively. When both types of measurements are available, they can be processed sequentially. The updated states and covariance matrix are then prepared for the next state prediction procedure.
V Simulation Results
In this section, we conduct simulation experiments to validate the superiority of WCI over classical CI in cooperative localization. We also compare the proposed DCL framework with the CCL framework to demonstrate its superiority in terms of robustness, as well as reduced computation and communication burdens.
Consider a 3D cooperative localization system where all agents initially remain static, with random positions and random attitudes (with roll and pitch set to zero). They then perform the following motion:
  • Accelerates at $0.4\,\text{m/s}^{2}$for 10 seconds.
  • Executes an 8-shaped trajectory (composed of two circular motions in opposite directions) for 26 seconds.
  • Elevates the pitch angle by 15 degrees over 1 second.
  • Moves at a constant velocity for 19 seconds.
  • Decreases the pitch angle by 15 degrees over 1 second.
  • Moves at a constant velocity for 13 seconds.
  • Performs a 180-degree clockwise circular turn over 20 seconds.
  • Moves at a constant velocity for 20 seconds.
  • Decelerates linearly to a hover over 10 seconds.
The total motion lasts 120 seconds, encompassing a variety of typical movements and sufficiently activating the observability of all state components. To prevent the agents’ states from diverging, we set up four anchors with known positions, located at $100,0]^{\mathsf{T}}\text{m}, [100,100,$$100]^{\mathsf{T}}\text{m}, [100,100,$$100,0]^{\mathsf{T}}\text{m}$ and $100,-100,100]^{\mathsf{T}}\text{m}$, respectively. The IMU parameters are set according to the ADIS16465 [30], which is a typical precise MEMS IMU. Regarding ranging techniques, we choose UWB, which can achieve centimeter-level ranging accuracy under line-of-sight conditions [31]. Unless otherwise specified, the remaining simulation parameters are set as shown in Table II. The trajectories of the agents and the positions of the anchors are depicted in Fig. 4.
We assume that agent 8 is connected to all four anchors, while agents 2-7 are connected to only two of the four anchors, which is common in practice due to communication range limitations or obstructions. However, agent 1 cannot connect to any anchors and can only correct its state by cooperating with other agents.
TABLE II: Default Simulation Parameters
Notation Value Description
$N_{u}$ 8 Number of agents
$f_{\text{uwb}}$ $5\,\text{Hz}$ Measurement frequency of UWB
$\sigma_{r}$ $0.1\,\text{m}$ Standard deviation of UWB’s measurement
$\boldsymbol{l}^{b}$ $0.1,0,0.1]^{\mathsf{T}}$m Lever arm from IMU to UWB in frame $\mathcal{B}$
$\sigma_{p_{0}}$ $0.3\,\text{m}$ Standard deviation of each agent’s initial position
$\sigma_{v_{0}}$ $0.1\,\text{m/s}$ Standard deviation of each agent’s initial velocity
$\sigma_{\phi_{0}}$ $3^{\circ}$ Standard deviation of each agent’s initial attitude
$T_{a}$ $5\,\text{s}$ Hyper-parameter of WCI
$N_{\text{sim}}$ 10 Total number of simulation runs
Figure 4: Anchor distributions and trajectories of the agents in the simulation experiments
We select four methods for comparison with the proposed WCI method.
  1. 1.
    NCL: No cooperative localization, where each agent corrects its state using only distance measurements from anchors. This method serves as a baseline to highlight the impact of inter-agent measurements and cooperation.
  2. 2.
    EKF: A CCL method where the computing center [^2] Unless otherwise specified, we assume that all IMU and distance measurements can be transmitted to the computing center without packet loss. maintains a large covariance matrix for all agents’ states and utilizes EKF for updates using all available measurements. Since it tracks the correlations between agents, it is considered the performance ceiling for cooperative localization.
  3. 3.
    CI-trace: A DCL method where each agent runs its own filter and uses trace-based CI for updates with measurements between agents.
  4. 4.
    CI-det: A DCL method where each agent runs its own filter and uses determinant-based CI for updates with measurements between agents.
We evaluate the state estimation results based on the state components of position, velocity, attitude, gyroscope bias, and accelerometer bias, respectively. The evaluation metrics used include root mean square error (RMSE), standard deviation (STD), and normalized estimation error squared (NEES). RMSE reflects the error between the estimated states and the true values, while STD indicates how well the nominal covariance matrix represents the uncertainty. NEES measures the relative magnitude of the estimation error compared to its covariance matrix, reflecting the consistency of the estimate [^3] Ideally, for a random vector of dimension $n$, the theoretical value of NEES is $n$; if NEES is significantly greater or smaller than $n$, the estimator is either overly pessimistic or optimistic. Let $\boldsymbol{y}$ denote the state component under consideration. At the time step $k$, the state estimate and true value are denoted as $\hat{\boldsymbol{y}}_{k}$ and $\boldsymbol{y}_{k}$, respectively, with the nominal covariance matrix given by $\boldsymbol{P}_{y,k}$. The formulas for these three metrics are as follows:
$\operatorname{RMSE}_{y,k}=\sqrt{\frac{1}{N_{\text{sim}}}\sum_{n=1}^{N_{\text{sim}}}(\hat{\boldsymbol{y}}_{k}-\boldsymbol{y}_{k})^{\mathsf{T}}(\hat{\boldsymbol{y}}_{k}-\boldsymbol{y}_{k})}$ (34)
$\operatorname{STD}_{y,k}=\sqrt{\frac{1}{N_{\text{sim}}}\sum_{n=1}^{N_{\text{sim}}}\operatorname{tr}(\boldsymbol{P}_{y,k})}$ (35)
$\operatorname{NEES}_{y,k}=\frac{1}{N_{\text{sim}}}\sum_{n=1}^{N_{\text{sim}}}(\hat{\boldsymbol{y}}_{k}-\boldsymbol{y}_{k})^{\mathsf{T}}\boldsymbol{P}_{y,k}^{-1}(\hat{\boldsymbol{y}}_{k}-\boldsymbol{y}_{k})$ (36)
where $N_{\text{sim}}$ is the total number of simulation runs.
V-A DCL Analysis
In this subsection, we analyze the performance of the proposed WCI-based DCL algorithm, comparing it with all benchmarks, particularly classical CI. The simulation settings are set to their default values. The RMSE and STD for all state components using different methods are shown in Fig. 5. The averaged RMSE, STD, and NEES over all time steps for all state components using different methods are presented in Table IIIb. The results for agent 1 and agents 2-8 are listed separately due to their different connections to anchors.
Figure 5: The RMSE and STD for all state components using different methods. (a) agent 1. (b) agent 2-8 (averaged).
TABLE III: RMSE, STD and NEES for Different Methods and State Components (Averaged over all time steps)
Method Position [m] Velocity [m/s] Attitude [deg] Gyro Bias [deg/h] Accel Bias $\boldsymbol{\mu}$g]
RMSE STD NEES RMSE STD NEES RMSE STD NEES RMSE STD NEES RMSE STD NEES
NCL 3982 3.0E+04 394.2 117.2 1.0E+04 7.702 30.96 3.0E+05 1.624 2495 1.4E+08 1.480 2425 1.4E+08 1.479
EKF 0.064 0.060 3.495 0.032 0.029 4.569 0.921 0.726 5.481 271.4 266.4 6.410 590.2 456.2 12.01
CI-trace 0.223 1.053 0.296 0.249 1.261 0.243 53.49 169.2 3.643 2.1E+04 6.7E+04 1.879 5.4E+04 1.5E+05 2.178
CI-det 0.458 1.660 0.250 0.120 0.695 0.129 2.692 29.36 0.186 499.6 4672 0.162 1208 1.5E+04 0.150
WCI 0.190 1.289 0.157 0.087 0.636 0.083 2.171 27.27 0.088 441.0 5022 0.080 1023 1.3E+04 0.075
(a) RMSE, STD and NEES of agent 1
Method Position [m] Velocity [m/s] Attitude [deg] Gyroscope Bias [deg/h] Accel Bias $\boldsymbol{\mu}$g]
RMSE STD NEES RMSE STD NEES RMSE STD NEES RMSE STD NEES RMSE STD NEES
NCL 8.394 1.614 8.0E+04 0.330 0.123 1249 1.350 0.909 36.85 312.2 301.7 27.52 962.3 673.3 117.0
EKF 0.053 0.049 3.511 0.028 0.025 4.419 0.946 0.681 5.304 272.4 256.7 6.001 536.5 434.2 8.736
CI-trace 0.144 0.338 2.491 0.124 0.234 4.117 21.19 22.68 8.275 4971 6906 4.922 1.1E+04 1.7E+04 3.322
CI-det 0.568 0.806 2.607 0.066 0.083 2.810 1.157 0.954 3.982 300.2 316.5 3.868 642.8 658.3 4.131
WCI 0.155 0.405 2.002 0.047 0.078 1.860 1.180 1.310 2.148 304.2 377.0 2.058 631.9 881.8 1.741
(b) Averaged RMSE, STD and NEES of agents 2-8
First, compared to the non-cooperative localization, all cooperative localization algorithms significantly improve state estimation, demonstrating that cooperation greatly enhances the usability of state estimation, especially position, under conditions of insufficient measurements with anchors. Among cooperative localization algorithms, it can be observed that the CCL EKF algorithm consistently provides the optimal estimates for all state components. In contrast, DCL algorithms, which do not track inter-agent correlations, show inferior state estimation performance compared to EKF. Furthermore, CI-based algorithms (including WCI) are inherently conservative estimation methods, with STD being higher than RMSE, which is an inherent characteristic of CI. When comparing agent 1 and agents 2-8, the primary difference lies in the availability of measurements from anchors. It is evident that nodes with anchor measurements integrated into the EKF update show some improvement in estimation accuracy, while their estimation conservativeness is significantly reduced. Additionally, when examining the estimation of different state components, it can be observed that the CI-trace method performs well in position estimation, with relatively small RMSE and the smallest STD among DCL algorithms. This is primarily due to the fact that its optimization objective is predominantly driven by position uncertainty. However, for other state components with smaller scales, such as attitude, gyroscope bias, and accelerometer bias, the CI-trace performs poorly. The attitude estimation error reaches several tens of degrees and the bias estimates also fail to serve as an inertial alignment correction. This has a detrimental impact on the agent: On the one hand, many tasks require not only precise position but also accurate attitude control; on the other hand, inaccurate attitude and bias estimates significantly impair the agent’s ability of autonomous navigation. When the agent fails to obtain sufficient distance measurements, the errors by only inertial navigation quickly accumulate and diverge. CI-det, on the other hand, achieves better estimation results for state components with smaller scales including attitude and IMU biases. However, it performs the worst in position estimation. In contrast, WCI achieves a balanced estimation across all states, yielding satisfactory results.
WCI and classical CI differ only in the state update phase. We then conduct an in-depth comparison of their performance during the state update phase. Agent 1 is selected as the subject of study, as its state update involves a single update method, unaffected by the interference of EKF updates based on distance measurements to anchors. Fig. 6 shows the change in the trace of the covariance matrix for each state component after each update, reflecting the variation in the uncertainty of the states. Since the trace changes exhibit significant variability and can be both positive and negative, a symmetric logarithmic scale is employed for the Y-axis (with non-equidistant intervals). We observe that, in contrast to EKF, which consistently reduces the uncertainty of each state component after each update, CI-based methods fail to achieve such consistency, especially for states with smaller scales, such as attitude, gyroscope bias, and accelerometer bias. For these states, it is evident that, in some cases, the updates lead to an increase in uncertainty, with the trace changes displaying an oscillatory pattern. However, compared to CI-trace, both CI-det and WCI better balance the optimization of states with varying scales. States with larger scales, such as position and velocity, show smaller changes in uncertainty after each update, while the oscillations of smaller scale states are reduced. This enables CI-det and WCI to achieve a more effective balance in optimizing uncertainty across different states, as illustrated in Fig. 5 and Table IIIb, thus preventing the severe divergence in uncertainty for smaller magnitude states observed in CI-trace.
Figure 6: The change in the trace of the covariance matrix for each state component after each update
We compared the nominal correlations imposed by the covariance matrices for the CI-trace, CI-det, and WCI methods, along with the actual correlations of the states. The average differences between the nominal and actual correlations are shown in Fig. 7. It is evident that, under all three methods, the actual correlations differ from the nominal ones, which is an inherent issue of CI that is difficult to avoid. However, in this experiment, compared to CI-det, both WCI and CI-trace exhibit lower correlation mismatches, particularly in the correlations related to position (the top three rows of the matrix), which dominate the state update. This explains why CI-det, while reflecting a better optimization balance between states of different scales, shows degradation in position estimation performance. In fact, the actual advantage of WCI lies in its flexibility, allowing the design of a weighting matrix $\boldsymbol{W}$ to balance different scale state estimations and address correlation mismatches in cooperative localization problems.
Figure 7: The average differences between the nominal and actual correlations
We also compared the impact of the hyperparameter $T_{a}$ chosen for the weighting matrix in WCI. The results are shown in Fig. 8, where we evaluate the performance of WCI with different weighting matrix by selecting different values of the hyperparameter $T_{a}$ in (30) and (32). Since CI-based methods are conservative estimators, STD and RMSE do not perfectly correspond to each other. WCI primarily optimizes based on STD. A larger value of $T_{a}$ indicates a greater emphasis on the long-term performance of the inertial navigation system during optimization, which places more weight on the uncertainty of states with higher-order relationships to propagation errors, such as attitude and IMU biases. This is reflected in the figure, where, as $T_{a}$ increases, the uncertainty in the position estimate gradually increases, while the uncertainty in the estimates of attitude and IMU biases decreases.
Figure 8: State estimation performance versus different values of $T_{a}$ for WCI
At the end of this subsection, we evaluate the advantages of the concurrent update in the proposed DCL algorithm. We compare it with the sequential update counterpart, where each agent updates its states immediately upon receiving one measurement. As shown in Fig. 9, EKF demonstrates little difference between the two schemes. However, for WCI, it is evident that the proposed concurrent update better aligns with the characteristics of CI fusion, exhibiting lower conservatism and superior estimation performance compared to sequential updates.
Figure 9: Comparison of update schemes: concurrent updates versus sequential updates
V-B Comparison Between DCL and CCL
In this subsection, we compare the DCL and CCL algorithms from two aspects. On one hand, we evaluate their robustness, focusing on the algorithms’ ability to adapt to scenarios involving communication packet loss and changes in swarm topology. This reflects whether the algorithm can remain effective and stable under complex and harsh conditions. On the other hand, we assess the computational and communication overhead of each method, which is a crucial factor influencing the scalability of the swarm and determines whether the algorithm can be applied to large-scale systems.
V-B1 Robustness
Figure 10: Position error comparison between DCL and CCL algorithms. After 30 seconds of motion, there is a 5% probability of data transmission failure from agent 1 to the data center. At 90 seconds, agent 1 goes offline and loses connection with both the computing center and other agents. It then rejoins the network at 110 seconds, restoring its connection with other agents.
The CCL algorithm requires all information to be gathered at the computing center. If packet loss or certain agent failure occurs, it not only halts the state estimation of the affected agent but also leads to incorrect correlation maintenance between agents. This, in turn, gradually affects the state estimation of other agents as well. We evaluate the robustness of the distributed and centralized algorithms under two conditions: (1) Communication packet loss, where data from certain nodes fails to reach the computing center, and (2) Changes in the network topology, where nodes go offline or join the system. It is assumed that, starting from 30 seconds, there is a 5% probability of data transmission failure from agent 1 to the data center. At 90 seconds, agent 1 goes offline, losing its connection with the computing center and other agents. Agent 1 then re-joins the network at 110 seconds, restoring its previous connections.
As shown in Fig. 10, we observe that after 30 seconds, not only does the position estimation error of agent 1 gradually increase, but it also skews the estimation results of other agents that have not experienced packet loss. At this point, the actual RMSE exceeds the estimated STD, indicating inconsistent estimates. In contrast, the DCL methods are completely unaffected by this issue.
During the 90-110s period, agent 1 is fully offline and disconnected with all anchors and other agents, relying solely on its own IMU outputs for state propagation. The state estimation error increases rapidly. However, methods like WCI and CI-det, which achieved balanced state estimation in earlier stages, show smaller attitude and IMU biases errors and uncertainties, leading to slower error accumulation during the autonomous navigation phase. Furthermore, when agent 1 reconnects, its state is not updated by the centralized algorithm during its offline period. As a result, when agent 1 re-joins, its differential distance measurements with other agent exhibit significant bias, which in turn skews the position estimates of the other agents. In contrast, the DCL algorithm allows the estimation of each agent to remain independent, making it more robust.
V-B2 Computation and Communication Overhead
To evaluate computational complexity, we compared the total processing time of the DCL algorithm and the CCL algorithm under different swarm sizes, as shown in Fig. 11. The total computational cost of the DCL algorithm increases linearly with the number of agents. When averaged on each agent, the computational overhead for its own state estimation remains almost constant as the swarm size increases, making it more suitable for large-scale swarms. In contrast, as the number of agents increases, the dimension of the covariance matrix in the CCL algorithm grows linearly, resulting in a cubic growth in computational cost (since the computational complexity of matrix multiplication and inversion involved in state prediction and update are all cubic in terms of matrix dimensions [32]). As a result, the processing time grows rapidly, making it difficult to scale to large swarms.
As for the communication load, the CCL algorithm requires high-frequency IMU data and measurement data to be collected at the computing center. Additionally, agents rely on messages from the computing center to obtain their state estimates. Some distributed implementations of CCL algorithms can also broadcast their state and state transition matrices to track the correlations between agents [33], which, although optimized to some extent, still require the transmission of large amounts of data. In contrast, our distributed approach only requires each agent to broadcast its position and attitude estimates, along with their variances, amounting to just a few dozen bytes of data, significantly reducing the communication burden of the system. Since this conclusion is evident, no separate simulation experiments are conducted for further analysis.
Figure 11: The total processing time of the DCL algorithm and the CCL algorithm under different swarm sizes
VI Conclusion
In this paper, we propose a WCI-based DCL algorithm designed for MAS operating in 3D scenarios. We identify three major drawbacks of CI in recursive filtering, which may pose risks to recursive filtering tasks, such as poor performance in 3D cooperative localization. We discover that a weighted mechanism, specifically WCI, has the potential to mitigate the inherent drawbacks of CI in recursive filtering tasks. Building on this, we introduce WCI into range-based DCL in 3D scenarios, designing a fusion strategy for multiple measurements and a weighting matrix selection strategy guided by the error propagation rule of INS for the cooperative localization task. Simulation results confirm that the proposed WCI algorithm outperforms classical CI in 3D cooperative localization tasks, and the proposed distributed framework achieves higher precision and tighter uncertainty compared to sequential updates. Additionally, the distributed framework demonstrates greater robustness and scalability, making it suitable for large-scale MAS. Future work will focus on developing dynamic weighting matrix selection strategies to adapt to the agents’ motion and cluster configuration changes, as well as applying more advanced distributed data fusion methods to improve the conservative nature of CI.
Appendix A Proof of Proposition 2
Reinhardt et al. studied the optimality of CI in [34]. We use its conclusion as a lemma:
Lemma 1.
For any strictly monotonically increasing cost function $J(\cdot)$ ($\boldsymbol{P}_{1}\succ\boldsymbol{P}_{2}\Rightarrow J(\boldsymbol{P}_{1})>J(\boldsymbol{P}_{2})$), CI provides the optimal bound in the fusion of two estimates under unknown correlations.
Based on Lemma 1, to prove Proposition 2, we should only prove that for any positive-definite weight matrix $\boldsymbol{W}$, $\operatorname{tr}(\boldsymbol{W}\boldsymbol{P})$ is a strictly monotonically increasing cost function, i.e., if and only if $\boldsymbol{W}$ is positive-definite, for any positive-definite matrices $\boldsymbol{P}_{1}\succ\boldsymbol{P}_{2}$, it holds that
$\operatorname{tr}(\boldsymbol{W}\boldsymbol{P}_{1})>\operatorname{tr}(\boldsymbol{W}\boldsymbol{P}_{2})\leftrightarrow\operatorname{tr}(\boldsymbol{W}(\boldsymbol{P}_{1}-\boldsymbol{P}_{2}))>0$ (37)
We prove Proposition 2 in two parts: sufficiency and necessity.
  • Sufficiency
Since $\boldsymbol{P}_{1}\succ\boldsymbol{P}_{2}$, $\boldsymbol{\Delta}=\boldsymbol{P}_{1}-\boldsymbol{P}_{2}$ is positive-definite. The positive-definite matrix $\boldsymbol{W}$ can be expressed as $\boldsymbol{W}=\boldsymbol{Q}^{\mathsf{T}}\boldsymbol{Q}$ for certain invertible $\boldsymbol{Q}$. Then
$\begin{split}\operatorname{tr}\left(\boldsymbol{W}(\boldsymbol{P}_{1}-\boldsymbol{P}_{2})\right)&=\operatorname{tr}\left(\boldsymbol{Q}^{\mathsf{T}}\boldsymbol{Q}\boldsymbol{\Delta}\right)\\ &=\operatorname{tr}\left(\boldsymbol{Q}\boldsymbol{\Delta}\boldsymbol{Q}^{\mathsf{T}}\right)\end{split}$ (38)
The second equation applies the commutative law of matrix trace [32]. Let $\boldsymbol{S}=\boldsymbol{Q}\boldsymbol{\Delta}\boldsymbol{Q}^{\mathsf{T}}$. Since $\boldsymbol{\Delta}$ is positive-definite and $\boldsymbol{Q}$ is invertible, $\boldsymbol{S}$ is positive-definite as well. Therefore, $\operatorname{tr}(\boldsymbol{S})=\operatorname{tr}(\boldsymbol{W}(\boldsymbol{P}_{1}-\boldsymbol{P}_{2}))>0$.
  • Necessity
If the weight matrix $\boldsymbol{W}$ is not positive-definite, then there must exist $\boldsymbol{x}\neq\boldsymbol{0}$ which satisfies $\boldsymbol{x}^{\mathsf{T}}\boldsymbol{W}\boldsymbol{x}\leq 0$. Let $\boldsymbol{\Delta}=\boldsymbol{P}_{1}-\boldsymbol{P}_{2}=\boldsymbol{x}\boldsymbol{x}^{\mathsf{T}}$ ($\boldsymbol{\Delta}$ is clearly positive-definite), then
$\operatorname{tr}(\boldsymbol{W}\boldsymbol{\Delta})=\operatorname{tr}(\boldsymbol{W}\boldsymbol{x}\boldsymbol{x}^{\mathsf{T}})=\boldsymbol{x}^{\mathsf{T}}\boldsymbol{W}\boldsymbol{x}\leq 0$ (39)
which is conflicted with the condition that $\operatorname{tr}(\boldsymbol{W}(\boldsymbol{P}_{1}-\boldsymbol{P}_{2}))>0$ for any positive-definite matrices $\boldsymbol{P}_{1}\succ\boldsymbol{P}_{2}$. So the assumption is not valid, and the weight matrix $\boldsymbol{W}$ must be positive-definite.