### v2

# Weighted Covariance Intersection for Range-based Distributed Cooperative Localization of Multi-Vehicle Systems

Abstract
Cooperative localization is considered a key solution for enabling autonomous navigation of multi-vehicle systems (MVS) in GNSS-denied environments. Among all solutions, distributed cooperative localization (DCL) has garnered widespread attention due to its robustness and scalability, making it well-suited for large-scale MVS. To address the challenge of untrackable state correlations between vehicles in a distributed framework, covariance intersection (CI) has been introduced as a means to fuse relative measurements under unknown correlations. However, existing studies treat CI merely as a plug-in method, applying traditional optimization criteria directly and focusing only on simple two-dimensional (2D) scenarios. When directly extended to three-dimensional (3D) scenarios with more complex state space (higher dimensions, additional state components, and significant disparities in scale and observability among state components), traditional methods fail to achieve balanced state estimation across all state components, leading to a significant degradation in the estimation accuracy of some state components. This highlights the need to design specialized mechanisms to improve the data fusion process. In this paper, we introduce a weighting mechanism, namely the weighted covariance intersection (WCI), to regulate the fusion process of CI. A concurrent fusion strategy for multiple distance measurements and a dedicated weighting matrix based on the error propagation rule of the inertial navigation system (INS) are developed for the data fusion process in DCL. Simulation results demonstrate that the proposed WCI significantly enhances cooperative localization performance compared to traditional CI, while the distributed approach outperforms the centralized approach in terms of robustness and scalability.
I Introduction
Autonomous vehicles, represented by unmanned aerial vehicles (UAVs) and unmanned ground vehicles (UGVs), have found wide applications in various fields such as environmental monitoring, disaster rescue, and military operations [1, 2, 3]. In recent years, multi-vehicle systems (MVS) have gained significant attention due to their high flexibility, collaboration capabilities, and robustness, allowing them to efficiently perform complex tasks [4, 5]. Accurate positioning services are the foundation for MVS to perform tasks. Currently, the Global Navigation Satellite System (GNSS) is one of the most widely used positioning methods in MVS [6]. Although GNSS provides all-day and all-weather positioning services in outdoor environments, its signals can be severely obstructed and become unavailable in indoor and underwater scenarios[7]. Local positioning systems (LPSs) require the prior deployment of infrastructure[8]. In contrast, cooperative localization, as an alternative solution, can effectively improve positioning accuracy and availability in GNSS-denied environments through vehicle collaboration[9, 10, 11].
Since the concept of cooperative localization was first introduced by [12], it has attracted widespread attention and made significant progress over the past three decades. Typically, cooperative localization involves relative measurements between vehicles, vehicle-to-vehicle communication (V2V) for information exchange, and the fusion of data of local and cooperative estimates[13], where well-localized vehicles can assist others to achieve better global localization performance [14]. Among various solutions, filter-based cooperative localization, which tracks vehicle poses (including both position and attitude) through recursive filtering, is particularly effective in enhancing autonomous navigation capabilities. Since each vehicle utilizes the state information of other vehicles to estimate its own state, their states become correlated, which must be accounted for during the filtering process when updating states. We refer to the state update process using measurements between vehicles as correlated update in later illustration. Depending on whether the computation occurs at a central unit or locally at each vehicle, cooperative localization can be divided into centralized cooperative localization (CCL) and distributed cooperative localization (DCL) [15]. Typically, the two schemes tackle the constantly changing state correlations between vehicles in different ways. CCL tracks the state correlations between vehicles with the requirement that all data be collected at a central processing unit, leading to significant communication overhead and poor scalability and robustness. In contrast, DCL typically neither requires nor is capable of tracking state correlations between vehicles, making it naturally more suitable for large-scale swarms and complex scenarios. However, this also presents the primary challenge of how to perform effective correlated update under unknown state correlations between vehicles.
Covariance intersection (CI) was initially proposed as a distributed data fusion method capable of producing a consistent and conservative estimate under unknown correlations between data sources [16], and has been extensively applied to recursive filtering tasks involving unknown correlations between measurements and estimated states, such as target tracking and cooperative perception [17, 18]. Since Arambel first introduced CI to cooperative localization [19], CI and its derivatives have been widely applied in cooperative vehicle localization with great success. Carrillo and Li used relative pose observations to obtain two estimates of the vehicle’s pose and fused them using CI and split covariance intersection (SCI), respectively [15, 20, 21]. Pierre extended CI-based correlated updates to partial observation scenarios, achieving DCL based solely on distance measurements [22]. Chang also utilized CI to develop a cooperative localization scheme that decouples measurement and communication, improving the robustness and scalability of distributed cooperative localization [23]. Fang considered the problem of cooperative localization under data sources of various error types and used the adaptive cubature split covariance intersection filter to estimate the vehicles’ states [24].
Despite the extensive research on CI-based cooperative localization, two major limitations remain. First, existing studies treat CI merely as a plug-in data fusion technique, directly applying traditional trace-based [21, 23] or determinant-based [15, 20, 22, 24] metrics of the fused covariance matrix, without considering how different optimization criteria affect the performance of cooperative localization tasks. Second, these studies are all designed for MVS moving in two-dimensional (2D) scenarios, where vehicles adopt unicycle model with a state space limited to three dimensions. There is still limited research on DCL for MVS, such as UAV swarms, operating in three-dimensional (3D) environments. Compared with 2D scenarios, the state space of vehicles is more complex, with higher dimensions, additional state components, and significant disparities in the scale and observability of state components. If existing approaches are directly applied to 3D vehicle cooperative localization, due to significant disparities in the scale and observability of state components, the CI fusion process under traditional CI criteria is dominated by position, and the uncertainty of state components with small scales and poor observability such as attitude will inflate greatly and their estimation performance degrades severely, as shown in Fig. 1. It is important to note that not only is attitude crucial in the operation of vehicles, but the estimation accuracy of small-scale states, such as attitude, also significantly impacts the long-term stability of localization. If the estimation errors of such states are large, the autonomous navigation capabilities of the vehicles will be severely affected, and the position error become highly susceptible to divergence, especially in situations without sufficient observation. We expect an improved CI optimization criterion that can balance and improve the estimation of all state components. As a result, specialized mechanisms must be designed to regulate and improve the data fusion process of CI in 3D cooperative localization task.
The weighting mechanism is widely employed to control the data fusion process [25]. Its primary advantage lies in its flexibility and controllability, enabling the adjustment of weights based on specific requirements to achieve optimal fusion results. By incorporating this weighting mechanism into CI, it’s promising to improve its performance in correlated data fusion, particularly in balancing state components with significant scale and observability disparities. Furthermore, the weighting mechanism provides a strategy to make the optimization criterion task-oriented. In practical applications, different state components can have varying impacts on the final result, and general mean squared error (MSE) optimization often fails to accurately reflect the actual contribution of the estimated states to the task at hand. In contrast, weighted mean squared error (WMSE) optimization offers clearer and more relevant significance in these scenarios [26]. Although the concept of weighted covariance intersection (WCI) was first introduced in [19], it primarily focused on enhancing the estimation performance of specific state subspaces and did not highlight the necessity of WCI in scenarios with significant disparities in the scale of state components. Furthermore, it lacked an interpretation of the optimization significance associated with WCI. These two limitations have, to some extent, restricted its application in recursive filtering tasks, such as cooperative localization. In contrast, in this paper, we carefully address these two challenges and successfully integrate WCI into DCL for 3D scenarios, achieving superior performance compared to traditional CI.
Figure 1: The influence of traditional CI criteria on 3D cooperative localization. For simplicity of illustration, we select the subspace of its whole state space with 2D position and attitude only.
In this paper, to achieve balanced estimation of different state components, we introduce WCI into range-based distributed cooperative localization in 3D scenarios. To improve the data fusion process, we develop a concurrent fusion strategy for multiple distance measurements that adapts to the characteristics of CI fusion and design a dedicated weighting matrix based on the error propagation rule of the inertial navigation system (INS). Extensive simulation experiments are conducted to validate the effectiveness and superiority of the proposed methods. The results show that WCI can effectively improve and balance state estimation in 3D cooperative localization tasks compared to traditional CI, and the distributed approach outperforms the centralized approach in terms of robustness and scalability.
The remainder of this paper is organized as follows. In Section II, we introduce the setup of the MVS under study, including the system and state composition, as well as range measurement models. In Section III, we introduce the overall distributed cooperative localization framework and focus our work on the correlated update process. Section IV elaborates on the proposed WCI-based correlated update process for DCL in 3D scenarios, including a concurrent fusion strategy for multiple distance measurements and a well-designed weighting matrix based on the error propagation rule of INS. In Section V, we present extensive simulations to validate the effectiveness of the proposed DCL method and demonstrate its superiority over CCL method. Finally, the conclusions are summarized in Section VII.
The main notations in the paper are summarized in Table I.
TABLE I: Notation List
Notation Description
lowercase $x$ scalar
bold lowercase $\boldsymbol{x}$ vector
bold uppercase $\boldsymbol{X}$ matrix
$\left\lVert\boldsymbol{x}\right\rVert$ Euclidean norm of a vector
$\mathbb{E}[\cdot]$ expectation operator
$\operatorname{tr}(\cdot)$ trace operator
$\operatorname{det}(\cdot)$ determinant operator
$\operatorname{diag}(\cdot)$ diagonal matrix with the entries inside
$(\cdot)^{\mathsf{T}}$ transpose operator
$\hat{(\cdot)}$ The term with a hat symbol above indicates that it is an estimate contaminated with noise.
$\tilde{(\cdot)}$ The term with a tilde symbol above indicates that it is a measurement contaminated with noise.
$(\cdot)\times$ The antisymmetric matrix constructed from the vector.
$\times$ multiplication cross operator
$\succeq$ Greater than or equal in the Löwner partial order. For example, $\boldsymbol{A}\succeq\boldsymbol{B}$ indicates that $\boldsymbol{A}-\boldsymbol{B}$ is a positive semidefinite matrix.
$\boldsymbol{I}_{M}$ $M\times M$ identity matrix
$\boldsymbol{0}_{M\times N}$ $M\times N$ zero vector or matrix
$\mathbb{R}^{M\times 1}$ $M\times 1$ real-valued vector
II System Setting
We consider a range-based cooperative localization system consisting of a MVS with $N_{v}$ vehicles moving in 3D space and $N_{a}$ fixed anchors with known positions, as shown in Fig. 2. Each vehicle is equipped with an onboard inertial measurement unit (IMU) to capture its self-motion. In addition, both vehicles and anchors are equipped with rigidly attached ranging modules. Through signal transceiving, each vehicle can measure the relative distances to other vehicles and anchors within its communication range [^1] In this study, anchors are assumed to function solely as signal transmitters, with measurements occurring exclusively on the vehicle side.. Furthermore, vehicles and anchors are equipped with communication modules for information exchange.
Since we use IMU as the proprioceptive sensor, the state of vehicle $i$ ($i=1,\ldots,N_{v}$) is defined as follows:
Figure 2: Schematic diagram of a range-based cooperative localization system. Due to communication range limitations or obstructions, vehicles may not be able to obtain sufficient distance measurements with anchors for state estimation. However, through vehicle collaboration, all vehicles can achieve accurate state estimation in such cases.
$\boldsymbol{x}_{i}=\left[\left(\prescript{I}{}{\boldsymbol{p}}_{i}^{n}\right)^{\mathsf{T}},\left(\prescript{I}{}{\boldsymbol{v}}_{i}^{n}\right)^{\mathsf{T}},\left(\boldsymbol{q}_{i}^{n}\right)^{\mathsf{T}},\boldsymbol{b\scriptstyle g}_{i}^{\mathsf{T}},\boldsymbol{b\scriptstyle a}_{i}^{\mathsf{T}}\right]^{\mathsf{T}}$ (1)
where $\prescript{I}{}{\boldsymbol{p}}_{i}^{n}$, $\prescript{I}{}{\boldsymbol{v}}_{i}^{n}$ and $\boldsymbol{q}_{i}^{n}$ represents the position, velocity and attitude (expressed in quaternion form) of the vehicle in the navigation frame $\mathcal{N}$, respectively. $\boldsymbol{b\scriptstyle g}_{i}$ and $\boldsymbol{b\scriptstyle a}_{i}$ represent the biases of the gyroscope and the accelerometer of the IMU, respectively. The superscript $I$ on the left indicates that the state corresponds to the position of the IMU on the vehicle. In IMU-based state estimation, system error propagation and state updates are typically performed using the error-state formulation rather than directly using the nominal state [27], which mitigates the redundancy inherent in quaternion-based attitude representation and addresses the challenges associated with nonlinearity. We define the error-state of vehicle $i$ as follows:
$\delta\boldsymbol{x}_{i}=\left[\left(\prescript{I}{}{\delta\boldsymbol{p}}_{i}^{n}\right)^{\mathsf{T}},\left(\prescript{I}{}{\delta\boldsymbol{v}}_{i}^{n}\right)^{\mathsf{T}},(\boldsymbol{\phi}^{n}_{i})^{\mathsf{T}},\left(\delta\boldsymbol{b\scriptstyle g}_{i}\right)^{\mathsf{T}},\left(\delta\boldsymbol{b\scriptstyle a}_{i}\right)^{\mathsf{T}}\right]^{\mathsf{T}}$ (2)
Here we adopt Phi-angle model $\boldsymbol{\phi}^{n}_{i}$ to represent the attitude error [28].
In general, the IMU and the ranging module are not co-located on the vehicle, and there is a lever arm between them. Assuming that each vehicle has an identical lever arm from the IMU to the ranging module, denoted as $\boldsymbol{l}^{b}$ in the vehicle’s body-fixed frame $\mathcal{B}$. The relationship between the IMU’s position and the ranging module’s position of vehicle $i$ can be expressed as follows:
$\prescript{R}{}{\boldsymbol{p}_{i}^{n}}=\prescript{I}{}{\boldsymbol{p}_{i}^{n}}+\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{l}^{b}$ (3)
where $\boldsymbol{C}_{b_{i}}^{n}$ denotes the rotation matrix that transforms a vector from frame $\mathcal{B}_{i}$ to frame $\mathcal{N}$. The relative distances between ranging modules can be expressed as
$d_{ij}^{\mathcal{V}}=\lVert\prescript{R}{}{\boldsymbol{p}_{i}^{n}}-\prescript{R}{}{\boldsymbol{p}_{j}^{n}}\rVert\qquad d_{im}^{\mathcal{A}}=\lVert\prescript{R}{}{\boldsymbol{p}_{i}^{n}}-\prescript{A}{}{\boldsymbol{p}_{m}^{n}}\rVert$ (4)
where $d_{ij}^{\mathcal{V}}$ is the distance between the ranging modules of vehicle $i$ and vehicle $j$, while $d_{im}^{\mathcal{A}}$ is the distance between the ranging modules of vehicle $i$ and anchor $m$ ($\prescript{A}{}{\boldsymbol{p}_{m}^{n}}$ is the position of anchor $m$ expressed in frame $\mathcal{N}$.). The superscripts $\mathcal{V}$ and $\mathcal{A}$ are used to distinguish the relative distances between two vehicles and those between an vehicle and an anchor. The distance measurements can be modeled as
${\tilde{d}_{ij}^{\mathcal{V}}}={{d}_{ij}^{\mathcal{V}}}+n_{d_{ij}^{\mathcal{V}}}\qquad{\tilde{d}_{im}^{\mathcal{A}}}={{d}_{im}^{\mathcal{A}}}+n_{d_{im}^{\mathcal{A}}}$ (5)
where $n_{d_{ij}^{\mathcal{V}}}$ and $n_{d_{im}^{\mathcal{A}}}$ are assumed to be i.i.d zero-mean Gaussian with variance $\sigma_{r}^{2}$.
III Distributed Cooperative Localization Framework
In this section, we introduce a range-based distributed cooperative localization framework for 3D scenarios. As illustrated in Fig. 3, we use the state estimation process of vehicle $i$ at time step $k$ as an example to explain the procedure of the general DCL framework.
In distributed cooperative localization, each vehicle only estimates its own state and maintains its own state covariance matrix. The state estimation of an vehicle can be divided into two phases: state prediction and state update. In the state prediction phase, the vehicle uses its onboard IMU outputs to predict state changes and adjust the covariance matrix for error propagation. In the state update phase, the vehicle uses distance measurements to other vehicles and anchors, along with information exchanged with other vehicles, to correct its state through data fusion. More specifically, distance measurements can be categorized into two types based on the object being measured: measurements between vehicles and anchors, and measurements between vehicles. Different data fusion methods are required for the two types of measurements.
III-A State Prediction Phase
The state prediction phase includes INS mechanism and error propagation, which has been extensively studied and are not the primary focus of our research. Therefore, we will not discuss these aspects here; readers are referred to [29, 27] for further details. It is worth mentioning that during the state prediction phase, states other than position, such as velocity, attitude, and IMU biases, affect the position estimation results. Therefore, for 3D cooperative localization, we must not only focus on the accuracy of position estimation but also ensure a balanced and satisfactory estimation of all state components.
Figure 3: Range-based distributed cooperative localization framework for 3D scenarios. Here $\{\boldsymbol{d}_{i}^{\mathcal{V}}\}$ and $\{\boldsymbol{d}_{i}^{\mathcal{A}}\}$ denote the sets of distance measurements from vehicle $i$ to other vehicles and anchors, respectively. $\text{info}_{i}$ and $\{\text{info}_{j}\}$ represent the information of vehicle $i$ that needs to transmitted to other vehicles, and the information received from other vehicles, respectively.
III-B State Update Phase
III-B1 Observation Equation
As for the state update phase, we begin with the observation equations for distance measurements. Suppose the predicted state of vehicle $i$ is defined as follows:
$\hat{\boldsymbol{x}}_{i}=\left[\left(\prescript{I}{}{\hat{\boldsymbol{p}}}_{i}^{n}\right)^{\mathsf{T}},\left(\prescript{I}{}{\hat{\boldsymbol{v}}}_{i}^{n}\right)^{\mathsf{T}},\left(\hat{\boldsymbol{q}}_{i}^{n}\right)^{\mathsf{T}},\hat{\boldsymbol{b\scriptstyle g}}_{i}^{\mathsf{T}},\hat{\boldsymbol{b\scriptstyle a}}_{i}^{\mathsf{T}}\right]^{\mathsf{T}}$
The predicted position of the ranging module can be written as:
$\begin{split}\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}&=\prescript{I}{}{\hat{\boldsymbol{p}}_{i}^{n}}+\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\\ &\approx(\prescript{I}{}{\boldsymbol{p}_{i}^{n}}+{\prescript{I}{}{\delta\boldsymbol{p}_{i}^{n}}})+\left[\boldsymbol{I}-({\boldsymbol{\phi}}_{i}^{n}\times)\right]\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{l}^{b}\\ &=\prescript{R}{}{\boldsymbol{p}_{i}^{n}}+{\prescript{I}{}{\delta\boldsymbol{p}_{i}^{n}}}+[(\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{l}^{b})\times]\boldsymbol{\phi}_{i}^{n}\end{split}$ (6)
Then, the predicted distance between vehicle $i$ and vehicle $j$, as well as the predicted distance between vehicle $i$ and anchor $m$ can be expressed as:
${\hat{d}_{ij}^{\mathcal{V}}}=$ $\left\lVert\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{R}{}{\hat{\boldsymbol{p}}_{j}^{n}}\right\rVert$ (7)
$\approx$ $\left\lVert\prescript{R}{}{\boldsymbol{p}_{i}^{n}}-\prescript{R}{}{\boldsymbol{p}_{j}^{n}}\right\rVert+(\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}})^{\mathsf{T}}\left(\prescript{I}{}{\delta{\boldsymbol{p}}_{i}^{n}}-\prescript{I}{}{\delta{\boldsymbol{p}}_{j}^{n}}\right)$
$+$ $(\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}})^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right]\boldsymbol{\phi}_{i}^{n}-(\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}})^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{j}}^{n}\boldsymbol{l}^{b}\right)\times\right]\boldsymbol{\phi}_{j}^{n}$
${\hat{d}_{im}^{\mathcal{A}}}=$ $\left\lVert\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{A}{}{\boldsymbol{p}}_{m}^{n}\right\rVert$
$\approx$ $\left\lVert\prescript{R}{}{\boldsymbol{p}_{i}^{n}}-\prescript{A}{}{\boldsymbol{p}}_{m}^{n}\right\rVert+(\hat{\boldsymbol{r}}_{im}^{\mathcal{A}})^{\mathsf{T}}\prescript{I}{}{\delta{\boldsymbol{p}}_{i}^{n}}$
$+$ $(\hat{\boldsymbol{r}}_{im}^{\mathcal{A}})^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right]\boldsymbol{\phi}_{i}^{n}$
where $\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}}$ and $\hat{\boldsymbol{r}}_{im}^{\mathcal{A}}$ denote the estimated unit direction vectors pointing from the positions of vehicle $j$’s and anchor $m$’s ranging modules to the position of vehicle $i$’s ranging module, respectively. Their explicit expressions are as follows:
$\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}}=\frac{\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{R}{}{\hat{\boldsymbol{p}}_{j}^{n}}}{\left\lVert\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{R}{}{\hat{\boldsymbol{p}}_{j}^{n}}\right\rVert}\qquad\hat{\boldsymbol{r}}_{im}^{\mathcal{A}}=\frac{\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{A}{}{{\boldsymbol{p}}_{m}^{n}}}{\left\lVert\prescript{R}{}{\hat{\boldsymbol{p}}_{i}^{n}}-\prescript{A}{}{{\boldsymbol{p}}_{m}^{n}}\right\rVert}$ (8)
Since we use the error-state formulation for error propagation and state updates, we apply differential measurements rather than the original measurements for state update. The differential measurements, i.e., the differences between the predicted distances and actual distance measurements, can be expressed as:
$\delta{d_{ij}^{\mathcal{V}}}$ $={\hat{d}_{ij}^{\mathcal{V}}}-{\tilde{d}_{ij}^{\mathcal{V}}}=\boldsymbol{H}_{ij}^{\mathcal{V}}\delta\boldsymbol{x}_{i}+\boldsymbol{G}_{ij}^{\mathcal{V}}\boldsymbol{v}_{ij}^{\mathcal{V}}$ (9)
$\delta{d_{im}^{\mathcal{A}}}$ $={\hat{d}_{im}^{\mathcal{A}}}-{\tilde{d}_{im}^{\mathcal{A}}}=\boldsymbol{H}_{ij}^{\mathcal{A}}\delta\boldsymbol{x}_{i}+v_{im}^{\mathcal{A}}$
Here, $\boldsymbol{v}_{ij}^{\mathcal{V}}$ and $v_{im}^{\mathcal{A}}$ are the noise of the differential measurements, where $v_{im}^{\mathcal{A}}=-n_{d_{im}^{\mathcal{A}}}$, and the expressions of $\boldsymbol{v}_{ij}^{\mathcal{V}}$ and its corresponding covariance matrix $\boldsymbol{R}_{v_{ij}}$ are given by
$\boldsymbol{v}_{ij}^{\mathcal{V}}=\left[\left(\prescript{I}{}{\delta{\boldsymbol{p}}_{j}^{n}}\right)^{\mathsf{T}},\left(\boldsymbol{\phi}_{j}^{n}\right)^{\mathsf{T}},n_{d_{ij}^{\mathcal{V}}}\right]^{\mathsf{T}}$ (10)
$\boldsymbol{R}_{ij}^{\mathcal{V}}=\begin{bmatrix}&\boldsymbol{P}_{{p}_{j}},&\boldsymbol{P}_{{p}_{j}{\phi}_{j}},&\boldsymbol{0}_{3\times 1}\\ &\boldsymbol{P}_{{p}_{j}{\phi}_{j}}^{\mathsf{T}},&\boldsymbol{P}_{{\phi}_{j}},&\boldsymbol{0}_{3\times 1}\\ &\boldsymbol{0}_{1\times 3},&\boldsymbol{0}_{1\times 3},&\sigma_{r}^{2}\end{bmatrix}$ (11)
Here $\boldsymbol{P}_{{p}_{j}}$, $\boldsymbol{P}_{{\phi}_{j}}$, $\boldsymbol{P}_{{p}_{j}{\phi}_{j}}$ are the submatrices of the state covariance matrix of vehicle $j$, corresponding to the covariance matrix of its position, attitude, and their cross-covariance, respectively. Additionally, the expressions of $\boldsymbol{H}_{ij}^{\mathcal{V}}$, $\boldsymbol{G}_{ij}^{\mathcal{V}}$ and $\boldsymbol{H}_{ij}^{\mathcal{A}}$ are given by
$\boldsymbol{H}_{ij}^{\mathcal{V}}$ $=\left[\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}}\right)^{\mathsf{T}},\boldsymbol{0}_{3\times 3},\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}}\right)^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right],\boldsymbol{0}_{3\times 6}\right]$ (12)
$\boldsymbol{G}_{ij}^{\mathcal{V}}$ $=\left[-\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}}\right)^{\mathsf{T}},-\left(\hat{\boldsymbol{r}}_{ij}^{\mathcal{V}}\right)^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right],-1\right]$
$\boldsymbol{H}_{im}^{\mathcal{A}}$ $=\left[\left(\hat{\boldsymbol{r}}_{im}^{\mathcal{A}}\right)^{\mathsf{T}},\boldsymbol{0}_{3\times 3},\left(\hat{\boldsymbol{r}}_{im}^{\mathcal{A}}\right)^{\mathsf{T}}\left[\left(\hat{\boldsymbol{C}}_{b_{i}}^{n}\boldsymbol{l}^{b}\right)\times\right],\boldsymbol{0}_{3\times 6}\right]$
Assume that vehicle $i$ obtains the distances to $M$ anchors and $N$ other vehicles for state updates at time step $k$. Without loss of generality, let the indices of the $M$ anchors and $N$ other vehicles be $1$ to $M$ and $1$ to $N$, respectively. Let the measurement vector formed by the differential distances to the $M$ anchors be denoted as $\boldsymbol{z}_{i}^{\mathcal{A}}=\left[\delta d_{i1}^{\mathcal{A}},\cdots,\delta d_{iM}^{\mathcal{A}}\right]^{\mathsf{T}}\in\mathbb{R}^{M\times 1}$, with the corresponding covariance matrix being $\boldsymbol{R}_{i}^{\mathcal{A}}=\sigma_{r}^{2}\boldsymbol{I}_{M}$; Let the differential distances to the $N$ other vehicles and their corresponding covariance matrices be denoted as $\left\{\delta d_{ij}^{\mathcal{V}},\boldsymbol{R}_{ij}^{\mathcal{V}}\right\}$, $j=1,\ldots,N$. The predicted state and covariance matrix of vehicle $i$ before the update are given by $\left\{\hat{\boldsymbol{x}}_{i,k|k-1},\boldsymbol{P}_{i,k|k-1}\right\}$.
III-B2 Independent Update
The differential distances between vehicle $i$ and anchors are independent of the vehicle’s state, and they can be directly used for state update via EKF. Since the distance measurements to the $M$ anchors are mutually independent, there is no difference between updating them sequentially as individual observations or combining them into a single observation vector for concurrent update. We choose the latter approach, and the corresponding update procedure can be written as:
$\boldsymbol{H}_{i}^{\mathcal{A}}=[{\boldsymbol{H}_{i1}^{\mathcal{A}}}^{\mathsf{T}},\cdots,{\boldsymbol{H}_{iM}^{\mathcal{A}}}^{\mathsf{T}}]^{\mathsf{T}}$
$\boldsymbol{K}=\boldsymbol{P}_{i,k|k-1}{\boldsymbol{H}_{i}^{\mathcal{A}}}^{\mathsf{T}}\left(\boldsymbol{H}_{i}^{\mathcal{A}}\boldsymbol{P}_{i,k|k-1}{\boldsymbol{H}_{i}^{\mathcal{A}}}^{\mathsf{T}}+\boldsymbol{R}_{i}^{\mathcal{A}}\right)^{-1}$
$\delta\boldsymbol{x}_{i,k}=\boldsymbol{K}\boldsymbol{z}_{i}^{\mathcal{A}}$ (13)
$\boldsymbol{P}_{i,k}=\left(\boldsymbol{I}-\boldsymbol{K}\boldsymbol{H}_{i}^{\mathcal{A}}\right)\boldsymbol{P}_{i,k|k-1}$
The estimated error-state $\delta\boldsymbol{x}_{i,k}$ can then be used to correct the estimated nominal state $\hat{\boldsymbol{x}}_{i,k|k-1}$, resulting in the updated nominal state $\hat{\boldsymbol{x}}_{i,k}$.
III-B3 Correlated Update
For distance measurements between vehicles, consider the measurement between vehicle $i$ and vehicle $j$ as an example, as shown in (9) and (10). The differential distance $\delta d_{ij}^{\mathcal{V}}$ is used for updating vehicle $i$’s state, and its noise includes the state estimation error of vehicle $j$. This leads to potential correlations between the differential distance measurement and vehicle $i$’s state. In a distributed framework, tracking the state correlation between vehicle $i$ and vehicle $j$ is extremely challenging or even infeasible. If the correlation is ignored and assumed to be independent, it could lead to the “data incest” problem [30], resulting in state estimation divergence. Therefore, our work focuses on this correlated update process, with the goal of achieving effective state update using relative distance measurements between vehicles under unknown correlations, ensuring accurate and consistent state estimation in 3D cooperative localization.
IV WCI-based Correlated Update
In this section, we present our methods for DCL in 3D scenarios. We begin by revisiting CI, including the version with weighting mechanism, WCI. Next, we propose a WCI-based approach for correlated updates in DCL, which accounts for the characteristics of CI fusion and the error propagation rule of INS, balancing the estimation of all state components with scale and observability disparities.
IV-A Covariance Intersection Revisit
Consider a state vector $\boldsymbol{x}$ with two unbiased and consistent estimates $\{\hat{\boldsymbol{{x}}}_{1},\boldsymbol{P}_{1}\}$ and $\{\hat{\boldsymbol{{x}}}_{2},\boldsymbol{P}_{2}\}$. Each estimate can be written as $\hat{\boldsymbol{{x}}}_{i}=\boldsymbol{x}+\delta{\boldsymbol{x}}_{i}$, $i=1,2$, where $\mathbb{E}[\delta{\boldsymbol{x}}_{i}]=\boldsymbol{0}$. We define the actual covariance matrices and the cross covariance matrix as
$\boldsymbol{P}_{i}^{*}=\mathbb{E}\left[\delta{\boldsymbol{x}}_{i}\delta{\boldsymbol{x}}_{i}^{\mathsf{T}}\right]\quad\boldsymbol{P}_{12}^{*}=\mathbb{E}\left[\delta{\boldsymbol{x}}_{1}\delta{\boldsymbol{x}}_{2}^{\mathsf{T}}\right]$ (14)
Since $\boldsymbol{P}_{i}^{*}$ are generally unknown, we name $\boldsymbol{P}_{1}$ and $\boldsymbol{P}_{2}$ as nominal covariance matrices. Consistency of each estimate then requires $\boldsymbol{P}_{i}\succeq\boldsymbol{P}_{i}^{*}$ [31].
CI produces an unbiased and consistent fused estimate $\{\hat{\boldsymbol{{x}}},\boldsymbol{P}\}$ without the knowledge of cross-correlation information as [16]
$\boldsymbol{P}$ $=\left(\omega\boldsymbol{P}_{1}^{-1}+(1-\omega)\boldsymbol{P}_{2}^{-1}\right)^{-1}$ (15)
$\hat{\boldsymbol{{x}}}$ $=\boldsymbol{\boldsymbol{P}}\left(\omega\boldsymbol{P}_{1}^{-1}\hat{\boldsymbol{x}}_{1}+(1-\omega)\boldsymbol{P}_{2}^{-1}\hat{\boldsymbol{x}}_{2}\right)$
where $\omega\in[0,1]$ is called the weight parameter. It has been proved that for any admissible $\omega$ and $\boldsymbol{P}_{12}^{*}$, the fused covariance matrix satisfies $\boldsymbol{P}\succeq\boldsymbol{P}^{*}$. The weight parameter $\omega$ is selected by minimizing a certain cost function $f(\boldsymbol{P};\omega)$:
$\omega^{*}=\arg\min_{\omega\in[0,1]}f(\boldsymbol{P};\omega)$ (16)
Typically, the cost function is chosen as the trace or determinant of the fused covariance matrix, i.e., $f(\boldsymbol{P})=\operatorname{tr}(\boldsymbol{P})$ or $\text{det}(\boldsymbol{P})$; we refer to them as the traditional criteria.
Furthermore, for partial-observation scenarios like $\boldsymbol{x}_{2}=\boldsymbol{H}\boldsymbol{x}$, where $\boldsymbol{H}$ is rank-deficient, and $\boldsymbol{P}_{1}^{-1}+\boldsymbol{H}^{\mathsf{T}}\boldsymbol{P}_{2}^{-1}\boldsymbol{H}$ is invertible, the CI estimate is given by [19]
$\boldsymbol{P}=\left(\omega\boldsymbol{P}_{1}^{-1}+(1-\omega)\boldsymbol{H}^{\mathsf{T}}\boldsymbol{P}_{2}^{-1}\boldsymbol{H}\right)^{-1}$ (17)
$\hat{\boldsymbol{{x}}}=\boldsymbol{\boldsymbol{P}}\left(\omega\boldsymbol{P}_{1}^{-1}\hat{\boldsymbol{x}}_{1}+(1-\omega)\boldsymbol{H}^{\mathsf{T}}\boldsymbol{P}_{2}^{-1}\hat{\boldsymbol{x}}_{2}\right)$
The traditional trace-based criterion is derived from the principle of minimizing MSE of the estimate—equivalently, the trace of the actual covariance matrix (cf. (18)). In practice, since the actual covariance matrix is unavailable, the trace of the nominal covariance matrix is minimized instead.
$\min\mathbb{E}[(\boldsymbol{x}-\boldsymbol{\hat{x}})^{\mathsf{T}}(\boldsymbol{x}-\boldsymbol{\hat{x}}))]=\operatorname{tr}(\boldsymbol{P}^{*})\rightarrow\min\operatorname{tr}(\boldsymbol{P})$ (18)
However, minimizing MSE alone is too coarse and lacks task specificity. To address this, Arambel et al. introduced relative weights to control the CI fusion process [19], termed WCI. However, they treated these weights merely as tuning knobs to improve the accuracy of a chosen component, without providing a theoretical rationale. We show in (19) that minimizing the weighted trace of the fused covariance matrix is mathematically equivalent to minimizing a WMSE over the entire state. By selecting the weighting matrix $\boldsymbol{W}$ to reflect task‐dependent priorities or target WMSE levels, WCI introduces task relevance and flexibility to the fusion criterion.
$\min\mathbb{E}[(\boldsymbol{x}-\boldsymbol{\hat{x}})^{\mathsf{T}}\boldsymbol{W}(\boldsymbol{x}-\boldsymbol{\hat{x}}))]=\operatorname{tr}(\boldsymbol{W}\boldsymbol{P}^{*})\rightarrow\min\operatorname{tr}(\boldsymbol{W}\boldsymbol{P})$ (19)
In many applications, all state components are crucial and must be accounted for in the data fusion process. However, when state components differ significantly in units or magnitudes, the CI fusion process under traditional criteria tends to be dominated by a particular state component, neglecting the others. To demonstrate this, we present a simple two-dimensional example. Let $\boldsymbol{P}_{1}=\begin{bmatrix}1&0\\ 0&0.01\end{bmatrix}$ and $\boldsymbol{P}_{2}=\begin{bmatrix}1.09&0\\ 0&0.001\end{bmatrix}$. Although $\boldsymbol{P}_{2}$ reduces the second component’s variance by 90% dramatically, its 9% increase in the first component drives the trace-minimization fusion (CI-trace) to favor $\boldsymbol{P}_{1}$. As shown in Fig. 4, the fused covariance matrix thus remains almost identical to $\boldsymbol{P}_{1}$, offering negligible improvement for the small-scale component. By contrast, determinant-based criterion (CI-det) favors $\boldsymbol{P}_{2}$, as the the determinant corresponds to the uncertainty volume in which components with small scales can influence the estimation. However, both criteria lack the ability to flexibly adjust and balance the estimation performance of different components in the presence of significant scale disparity. By selecting an appropriate weighting matrix and applying WCI, it becomes possible to effectively balance the fusion process and fully leverage the information from both sources.
Figure 4: CI fusion under scale disparities (axes scaled differently for better demonstration). The covariance ellipse of CI-trace almost overlaps that of $\boldsymbol{P}_{1}$, indicating that the fusion is dominated by the larger-scale state component. As a result, the fused estimate closely follows the lower-uncertainty estimate along that dimension. While CI-det can account for smaller-scale state component, it still fails to achieve a proper balance between two components. In contrast, WCI effectively balances the fusion process and makes full use of the information from both sources.
IV-B Concurrent Update Strategy
From this subsection, continuing from the distributed cooperative localization framework introduced in Section III, we present the WCI-based correlated update process. Using vehicle $i$ as an example, assume it has obtained differential distances to the $N$ other vehicles and their corresponding covariance matrices be denoted as $\left\{\delta d_{ij}^{\mathcal{V}},\boldsymbol{R}_{ij}^{\mathcal{V}}\right\}$, $j=1,\ldots,N$, whose explicit expressions are given in (9) and (11), respectively. Since the noise of the differential distance measurements contains estimation errors of other vehicles’ states, which are correlated with the state of vehicle $i$ and these correlations are unknown, we can use CI, or its enhanced version WCI, introduced in Section IV-A, to perform the correlated update process.
Unlike EKF, where sequential and concurrent updates of multiple observations yield the same result, CI does not provide the same guarantee. In fact, CI ensures that the fusion of two data sources provides an optimal conservative upper bound, but it cannot guarantee the same for the fusion of multiple data sources [32]. Sequential updates of multiple observations lead to additional performance degradation and result in a more conservative upper bound. Therefore, for multiple distance measurements between vehicles, we combine them into a single observation vector for state updates.
$\boldsymbol{z}_{i}^{\mathcal{V}}=\left[\delta{d_{i1}^{\mathcal{V}}},\cdots,\delta{d_{iN}^{\mathcal{V}}}\right]^{\mathsf{T}}=\boldsymbol{H}_{i}^{\mathcal{V}}\delta\boldsymbol{x}^{i}+\boldsymbol{G}_{i}^{\mathcal{V}}\boldsymbol{v}_{i}^{\mathcal{V}}$ (20)
where the explicit expressions of $\boldsymbol{H}_{i}^{\mathcal{V}}$ and $\boldsymbol{G}_{i}^{\mathcal{V}}$ are given by:
$\boldsymbol{H}_{i}^{\mathcal{V}}$ $=[{\boldsymbol{H}_{i1}^{\mathcal{V}}}^{\mathsf{T}},\cdots,{\boldsymbol{H}_{iN}^{\mathcal{V}}}^{\mathsf{T}}]^{\mathsf{T}}$ (21)
$\boldsymbol{G}_{i}^{\mathcal{V}}$ $=[{\boldsymbol{G}_{i1}^{\mathcal{V}}}^{\mathsf{T}},\cdots,{\boldsymbol{G}_{iN}^{\mathcal{V}}}^{\mathsf{T}}]^{\mathsf{T}}$
The combined noise vector and its actual covariance matrix can be written as:
$\boldsymbol{v}_{i}^{\mathcal{V}}$ $=\left[{\boldsymbol{v}_{i1}^{\mathcal{V}}}^{\mathsf{T}},\cdots,{\boldsymbol{v}_{iN}^{\mathcal{V}}}^{\mathsf{T}}\right]^{\mathsf{T}}$ (22)
$\boldsymbol{R}_{i}^{\mathcal{V}}$ $=\begin{bmatrix}\boldsymbol{R}_{i1}^{\mathcal{V}}&\cdots&\boldsymbol{C}_{i,1N}^{\mathcal{V}}\\ \vdots&\ddots&\vdots\\ {\boldsymbol{C}_{i,1N}^{\mathcal{V}}}^{\mathsf{T}}&\cdots&\boldsymbol{R}_{iN}^{\mathcal{V}}\end{bmatrix}$
Here, $\boldsymbol{C}_{i,1N}^{\mathcal{V}}$ denote the cross-covariance matrix between $\boldsymbol{v}_{i1}^{\mathcal{V}}$ and $\boldsymbol{v}_{iN}^{\mathcal{V}}$, which is generally non-zero since the states of vehicle 1 and vehicle $N$ are potentially correlated. However, $\boldsymbol{C}_{i,1N}^{\mathcal{V}}$ is not unavailable in practical distributed systems, as we do not track the correlations between vehicles. An alternative approach is to appropriately inflate the noise covariance matrix while ensuring that the nominal covariance matrix remains greater than the actual covariance matrix, thus maintaining consistency. We focus on the diagonal elements of $\boldsymbol{R}_{i}^{\mathcal{V}}$ and multiply the vehicle state error variance terms by 2 to obtain the nominal noise covariance matrix $\boldsymbol{R}_{i}^{\mathcal{V}^{\prime}}$, as follows:
$\boldsymbol{R}_{i}^{\mathcal{V}^{\prime}}=\begin{bmatrix}2\boldsymbol{P}_{p_{1}}^{d}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}&\cdots&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{3\times 3}&2\boldsymbol{P}_{\phi_{1}}^{d}&\boldsymbol{0}_{3\times 1}&\cdots&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&\sigma_{r}^{2}&\cdots&\boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}&\cdots&2\boldsymbol{P}_{p_{N}}^{d}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}&\cdots&\boldsymbol{0}_{3\times 3}&2\boldsymbol{P}_{\phi_{N}}^{d}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&0&\cdots&\boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&\sigma_{r}^{2}\end{bmatrix}$ (23)
where
$\boldsymbol{P}_{p_{j}}^{d}=\operatorname{diag}(\left[\sigma_{p_{j,x}}^{2},\sigma_{p_{j,y}}^{2},\sigma_{p_{j,z}}^{2}\right]^{\mathsf{T}})$ (24)
$\boldsymbol{P}_{\phi_{j}}^{d}=\operatorname{diag}(\left[\sigma_{\phi_{j,x}}^{2},\sigma_{\phi_{j,y}}^{2},\sigma_{\phi_{j,z}}^{2}\right]^{\mathsf{T}})$
($j=1,\ldots,N$) are the diagonal matrices consisting of the error variances of vehicle $j$’s position and attitude. It is easy to observe that $\boldsymbol{R}_{i}^{\mathcal{V}^{\prime}}\succeq\boldsymbol{R}_{i}^{\mathcal{V}}$, meaning that the nominal covariance matrix $\boldsymbol{R}_{i}^{\mathcal{V}^{\prime}}$ still provides a consistent estimate of the differential measurement vector. In addition, the nominal covariance matrix is free from the correlation items, thereby eliminating the need to track correlations and reducing the communication burden.
With the differential measurements between vehicles $\left\{\boldsymbol{z}_{i}^{\mathcal{V}},\boldsymbol{R}_{i}^{\mathcal{V}^{\prime}}\right\}$, we can apply CI for state updates. A KF-style CI formulation for partial observations is provided in [20] as follows:
$\boldsymbol{P}_{1}=\frac{1}{\omega}\boldsymbol{P}_{i,k|k-1}$
$\boldsymbol{P}_{2}=\frac{1}{1-\omega}\boldsymbol{R}_{i}^{\mathcal{V}^{\prime}}$
$\boldsymbol{K}=\boldsymbol{P}_{1}{\boldsymbol{H}_{i}^{\mathcal{V}}}^{\mathsf{T}}\left({\boldsymbol{H}_{i}^{\mathcal{V}}}\boldsymbol{P}_{1}{\boldsymbol{H}_{i}^{\mathcal{V}}}^{\mathsf{T}}+\boldsymbol{P}_{2}\right)^{-1}$ (25)
$\delta{\boldsymbol{x}}_{i,k}=\boldsymbol{K}\boldsymbol{z}_{i}^{\mathcal{V}}$
$\boldsymbol{P}_{i,k}=(\boldsymbol{I}-\boldsymbol{K}{\boldsymbol{H}_{i}^{\mathcal{V}}})\boldsymbol{P}_{1}$
Here, $\omega$ is the weight parameter, which is determined by a specific optimization criterion in (16).
In the context of our problem, the state vector includes components with large scale and high observability, such as position, as well as components with small scale and low observability, such as attitude and IMU biases. Traditional criteria, such as CI-trace, typically reduce the uncertainty of large scale and highly observable components like position, but at the cost of significantly increasing the uncertainty of other state components. This imbalance ultimately degrades the overall estimation performance. In contrast, CI-det focuses more on optimizing the smaller-scale variables, giving more attention to those components with low scale. As a result, both the traditional criterion of minimizing the trace or the determinant of the fused covariance matrix perform poorly in the 3D cooperative localization task. Instead, we choose WCI as presented in (19), with the primary challenge lying in the selection of the weighting matrix.
IV-C Weighting Matrix Selection
$\boldsymbol{S}=\begin{bmatrix}\boldsymbol{I}_{3},\ \boldsymbol{I}_{3}\cdot T_{a},\ \frac{1}{2}\left[\left(\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{f}_{i}^{b}\right)\times\right]T_{a}^{2},\ -\frac{1}{6}\left[\left(\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{f}_{i}^{b}\right)\times\right]\boldsymbol{C}_{b_{i}}^{n}T_{a}^{3},\ \frac{1}{2}\boldsymbol{C}_{b_{i}}^{n}T_{a}^{2}\end{bmatrix}$ (30)
$\mathbb{E}\left[\left(\prescript{I}{}{\delta{\boldsymbol{p}}}_{i,T_{a}}^{n}\right)^{\mathsf{T}}\left(\prescript{I}{}{\delta{\boldsymbol{p}}}_{i,T_{a}}^{n}\right)\right]=\operatorname{tr}\left(\mathbb{E}\left[\left(\prescript{I}{}{\delta{\boldsymbol{p}}}_{i,T_{a}}^{n}\right)\left(\prescript{I}{}{\delta{\boldsymbol{p}}}_{i,T_{a}}^{n}\right)^{\mathsf{T}}\right]\right)=\operatorname{tr}\left(\boldsymbol{S}\mathbb{E}\left[\delta\boldsymbol{x}_{i,0}\delta\boldsymbol{x}_{i,0}^{\mathsf{T}}\right]\boldsymbol{S}^{\mathsf{T}}\right)=\operatorname{tr}\left(\boldsymbol{S}^{\mathsf{T}}\boldsymbol{S}\boldsymbol{P}_{i,0}\right)$ (31)
Inertial navigation serves as the primary method for state propagation in the vehicle state estimation procedure, while distance measurements are used as a supplementary tool for state correction. If the error in inertial navigation can be minimized as much as possible, the overall state estimation error in cooperative localization will decrease, thereby enhancing the vehicle’s autonomous navigation ability even in scenarios without sufficient distance measurements. In summary, the optimization goal in the state update phase can be framed as minimizing the inertial navigation error. Additionally, inertial navigation error is related to all state components, and therefore, optimizing these errors can provide guidance for achieving a balanced ratio between different state components.
We consider the propagation law of inertial navigation errors. Generally, vehicles are equipped with low cost micro-electromechanical system (MEMS) sensors, so we do not account for factors such as the earth’s rotation, the variation of frame $\mathcal{N}$, or changes in gravity. As a result, a simplified error state model can be expressed as [28]:
$\left\{\begin{aligned} \prescript{I}{}{\delta{\dot{\boldsymbol{p}}}_{i}^{n}}&=\prescript{I}{}{\delta{\boldsymbol{v}}_{i}^{n}}\\ \prescript{I}{}{\delta{\dot{\boldsymbol{v}}}_{i}^{n}}&=\boldsymbol{f}_{i}^{n}\times\boldsymbol{\phi}_{i}^{n}+\boldsymbol{C}_{b_{i}}^{n}\delta\boldsymbol{f}_{i}^{b}\\ \dot{\boldsymbol{\phi}}_{i}^{n}&=-\boldsymbol{C}_{b_{i}}^{n}\delta\boldsymbol{\omega}_{i}^{b}\end{aligned}\right$. (26)
where $\boldsymbol{f}_{i}^{n}=\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{f}_{i}^{b}$ is the projection of the specific force $\boldsymbol{f}_{i}^{b}$ into frame $\mathcal{N}$. $\delta\boldsymbol{f}_{i}^{b}$ and $\delta\boldsymbol{\omega}_{i}^{b}$ represent the biases in the IMU outputs, specifically the specific force and angular velocity, with respect to their actual values. These biases are modeled as:
$\delta\boldsymbol{f}_{i}^{b}=\boldsymbol{b\scriptstyle a}_{i}+\boldsymbol{w}_{a}$ (28)
$\delta\boldsymbol{\omega}_{i}^{b}=\boldsymbol{b\scriptstyle g}_{i}+\boldsymbol{w}_{g}$
where $\boldsymbol{w}_{a}$ and $\boldsymbol{w}_{g}$ are zero-mean random noise. The error state model indicates that the accumulation of inertial navigation position errors is related to the initial position error by a zero-order relationship, which remains constant over time. It is related to the initial velocity error by a first-order relationship, which increases linearly with time. The position error is related to the initial attitude and acceleration errors by a second-order relationship, which increases quadratically with time, and it is related to the initial gyroscope bias error by a third-order relationship, which increases cubically with time[29].
We make a simple assumption that inertial navigation position error accumulates similarly to the case where vehicle $i$ moves with constant velocity in a straight line over a time period of $T_{a}$. The inertial navigation position error at time $T_{a}$ can then be approximately written as:
$\begin{split}\prescript{I}{}{\delta{\boldsymbol{p}}}_{i,T_{a}}^{n}=\ &\prescript{I}{}{\delta{\boldsymbol{p}}}_{i,0}^{n}+\prescript{I}{}{\delta{\boldsymbol{v}}}_{i,0}^{n}T_{a}+\frac{1}{2}\left[\left(\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{f}^{b}\right)\times\right]\boldsymbol{\phi}_{i}^{n}T_{a}^{2}+\\ &\frac{1}{2}\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{b\scriptstyle a}_{i}T_{a}^{2}-\frac{1}{6}\left[\left(\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{f}^{b}\right)\times\right]\boldsymbol{C}_{b_{i}}^{n}\boldsymbol{b\scriptstyle g}_{i}T_{a}^{3}\\ =\ &\boldsymbol{S}\delta\boldsymbol{x}_{i,0}\end{split}$ (29)
where the explicit expression for $\boldsymbol{S}$ is given by (30). We aim to minimize the mean squared error of $\prescript{I}{}{{\boldsymbol{p}}}_{i,T_{a}}^{n}$, which corresponds to minimizing the trace of the product of $\boldsymbol{S}^{\mathsf{T}}\boldsymbol{S}$ and initial error-state covariance matrix $\boldsymbol{P}_{i,0}$ as illustrated in (31).
It’s interesting to notice that if we assign the weighting matrix $\boldsymbol{W}$ as
$\boldsymbol{W}=\boldsymbol{S}^{\mathsf{T}}\boldsymbol{S}$ (32)
The optimization goal in (31) is the same as that in (19). We can finally determine the weight parameter $\omega$ in (25) by
$\omega^{*}=\arg\min_{\omega\in[0,1]}\operatorname{tr}(\boldsymbol{W}_{i,k}\boldsymbol{P}_{i,k})$ (33)
where $\boldsymbol{W}_{i,k}$ is determined by (30) and (32) with predicted state estimates and IMU outputs at time step $k$. The estimated error-state $\delta\boldsymbol{x}_{i,k}$ can be used to correct nominal state estimates as well. The entire procedures of the WCI-based correlated update are summarized in Algorithm.
Algorithm WCI-based Correlated Update Method
  • State estimate after prediction $\{\hat{\boldsymbol{x}}_{i,k|k-1},\boldsymbol{P}_{i,k|k-1}\}$
  • IMU measurement $\tilde{\boldsymbol{f}}_{k,i}^{b}$
  • Distance measurements with other vehicles $\{d_{ij}^{\mathcal{V}}\}$
  • Other vehicles’ position and attitude estimates $\prescript{\mathcal{I}}{}{\hat{\boldsymbol{p}}_{j,k|k-1}^{n}}$ and $\hat{\boldsymbol{q}}_{j,k|k-1}^{n}$, along with their nominal variances $\{\sigma_{p_{j,x}}^{2},\sigma_{p_{j,y}}^{2}\sigma_{p_{j,z}}^{2}\}$ and $\{\sigma_{\phi_{j,x}}^{2},\sigma_{\phi_{j,y}}^{2}\sigma_{\phi_{j,z}}^{2}\}$.
State estimate after update $\{\hat{\boldsymbol{x}}_{i,k},\boldsymbol{P}_{i,k}\}$
Calculate differential measurements with the inflated nominal covariance matrices $\{\boldsymbol{z}_{i}^{\mathcal{V}},\boldsymbol{R}_{i}^{\mathcal{V}^{\prime}}\}$ according to (17), (20), (23) and (24).
Determine the weighting matrix $\boldsymbol{W}_{i,k}$ by (30) and (32), replacing ${\boldsymbol{f}}_{k,i}^{b}$ with $\tilde{\boldsymbol{f}}_{k,i}^{b}$.
Execute data fusion using WCI by (25) and (33).
Perform error feedback to correct the predicted state estimate.
Furthermore, we have described the independent and correlated update procedures for distance measurements between vehicles and anchors, and between vehicles, using EKF and WCI, respectively. When both types of measurements are available, they can be processed sequentially. The updated states and covariance matrix are then prepared for the next state prediction procedure.
V Simulation Results
In this section, we conduct simulation experiments to validate the superiority of WCI over traditional CI in 3D cooperative localization. We also compare the proposed DCL framework with the CCL framework to demonstrate its superiority in terms of robustness, as well as reduced computation and communication burdens.
Consider a 3D cooperative localization system where all vehicles initially remain static, with random positions and random attitudes (with roll and pitch set to zero). They then perform the following motion:
  • Accelerates at $0.4\,\text{m/s}^{2}$for 10 seconds.
  • Executes an 8-shaped trajectory (composed of two circular motions in opposite directions) for 26 seconds.
  • Elevates the pitch angle by 15 degrees over 1 second.
  • Moves at a constant velocity for 19 seconds.
  • Decreases the pitch angle by 15 degrees over 1 second.
  • Moves at a constant velocity for 13 seconds.
  • Performs a 180-degree clockwise circular turn over 20 seconds.
  • Moves at a constant velocity for 20 seconds.
  • Decelerates linearly to a hover over 10 seconds.
The total motion lasts 120 seconds, encompassing a variety of typical movements and sufficiently activating the observability of all state components. To prevent the vehicles’ states from diverging, we set up four anchors with known positions, located at $100,0]^{\mathsf{T}}\text{m}, [100,100,$$100]^{\mathsf{T}}\text{m}, [100,100,$$100,0]^{\mathsf{T}}\text{m}$ and $100,-100,100]^{\mathsf{T}}\text{m}$, respectively. The IMU parameters are set according to the ADIS16465 [33], which is a typical precise MEMS IMU. Regarding ranging techniques, we choose UWB, which can achieve centimeter-level ranging accuracy under line-of-sight conditions [34]. Unless otherwise specified, the remaining simulation parameters are set as shown in Table II. The trajectories of the vehicles and the positions of the anchors are depicted in Fig. 5.
We assume that vehicle 8 is connected to all four anchors, while vehicles 2-7 are connected to only two of the four anchors, which is common in practice due to communication range limitations or obstructions. However, vehicle 1 cannot connect to any anchors and can only correct its state by cooperating with other vehicles.
TABLE II: Default Simulation Parameters
Notation Value Description
$N_{v}$ 8 Number of vehicles
$f_{\text{uwb}}$ $5\,\text{Hz}$ Measurement frequency of UWB
$\sigma_{r}$ $0.1\,\text{m}$ Standard deviation of UWB’s measurement
$\boldsymbol{l}^{b}$ $0.1,0,0.1]^{\mathsf{T}}$m Lever arm from IMU to UWB in frame $\mathcal{B}$
$\sigma_{p_{0}}$ $0.3\,\text{m}$ Standard deviation of each vehicle’s initial position
$\sigma_{v_{0}}$ $0.1\,\text{m/s}$ Standard deviation of each vehicle’s initial velocity
$\sigma_{\phi_{0}}$ $3^{\circ}$ Standard deviation of each vehicle’s initial attitude
$T_{a}$ $5\,\text{s}$ Hyper-parameter of WCI
$N_{\text{sim}}$ 10 Total number of simulation runs
Figure 5: Anchor distributions and trajectories of the vehicles in the simulation experiments
We select four methods for comparison with the proposed WCI method.
  1. 1.
    NCL: No cooperative localization, where each vehicle corrects its state using only distance measurements from anchors. This method serves as a baseline to highlight the impact of inter-vehicle measurements and cooperation.
  2. 2.
    CCL EKF: A CCL method where the computing center [^2] Unless otherwise specified, we assume that all IMU and distance measurements can be transmitted to the computing center without packet loss. maintains a large covariance matrix for all vehicles’ states and utilizes EKF for updates using all available measurements. Since it tracks the correlations between vehicles, it is considered the performance ceiling for cooperative localization.
  3. 3.
    DCL CI-trace: A DCL method where each vehicle runs its own filter and uses trace-based CI for updates with measurements between vehicles.
  4. 4.
    DCL CI-det: A DCL method where each vehicle runs its own filter and uses determinant-based CI for updates with measurements between vehicles.
We evaluate the state estimation results based on the state components of position, velocity, attitude, gyroscope bias, and accelerometer bias, respectively. The evaluation metrics used include root-mean-squared-error (RMSE), root-mean-trace-error (RMTE) [23], and normalized estimation error squared (NEES). RMSE reflects the error between the estimated states and the true values, while RMTE indicates how well the nominal covariance matrix represents the uncertainty. NEES measures the relative magnitude of the estimation error compared to its covariance matrix, reflecting the consistency of the estimate [^3] Ideally, for a random vector of dimension $n$, the theoretical value of NEES is $n$; if NEES is significantly greater or smaller than $n$, the estimator is either overly pessimistic or optimistic. Let $\boldsymbol{y}$ denote the state component under consideration. At the time step $k$, the state estimate and true value are denoted as $\hat{\boldsymbol{y}}_{k}$ and $\boldsymbol{y}_{k}$, respectively, with the nominal covariance matrix given by $\boldsymbol{P}_{y,k}$. The formulas for these three metrics are as follows:
$\operatorname{RMSE}_{y,k}=\sqrt{\frac{1}{N_{\text{sim}}}\sum_{n=1}^{N_{\text{sim}}}(\hat{\boldsymbol{y}}_{k}-\boldsymbol{y}_{k})^{\mathsf{T}}(\hat{\boldsymbol{y}}_{k}-\boldsymbol{y}_{k})}$ (34)
$\operatorname{RMTE}_{y,k}=\sqrt{\frac{1}{N_{\text{sim}}}\sum_{n=1}^{N_{\text{sim}}}\operatorname{tr}(\boldsymbol{P}_{y,k})}$ (35)
$\operatorname{NEES}_{y,k}=\frac{1}{N_{\text{sim}}}\sum_{n=1}^{N_{\text{sim}}}(\hat{\boldsymbol{y}}_{k}-\boldsymbol{y}_{k})^{\mathsf{T}}\boldsymbol{P}_{y,k}^{-1}(\hat{\boldsymbol{y}}_{k}-\boldsymbol{y}_{k})$ (36)
where $N_{\text{sim}}$ is the total number of simulation runs.
V-A DCL Analysis
In this subsection, we analyze the performance of the proposed WCI-based DCL algorithm, comparing it with all benchmarks, with a particular focus on traditional CI. The simulation parameters are set to their default values. Due to different connectivity with anchors, we select three representative vehicles, indexed as 1, 3, and 6, for analysis. The RMSE and RMTE for all state components using different methods are shown in Fig. 6. The averaged RMSE, RMTE, and NEES over all time steps for all state components using different methods are presented in Table IIIc.
Figure 6: The RMSE and RMTE for all state components using different methods. (a) vehicle 1. (b) vehicle 3. (c) vehicle 6
First, compared to the non-cooperative localization, all cooperative localization algorithms significantly improve state estimation, demonstrating that cooperation greatly enhances the usability of state estimation, especially position, under conditions of insufficient measurements with anchors. Among cooperative localization algorithms, it can be observed that the CCL EKF algorithm consistently provides the optimal estimates for all state components. In contrast, DCL algorithms, which do not track inter-vehicle correlations, show inferior state estimation performance compared to CCL EKF. Furthermore, CI-based algorithms (including WCI) are inherently conservative estimation methods, with RMTE being higher than RMSE, which is an inherent characteristic of CI. When comparing vehicle 1 and vehicles 3,6, the primary difference lies in the availability of measurements from anchors. It is evident that nodes with anchor measurements integrated into the EKF update generally show some improvement in estimation accuracy, while their estimation conservativeness is significantly reduced. Additionally, when examining the estimation of different state components, it can be observed that the CI-trace method performs well in position estimation, with relatively small RMSE and the smallest RMTE among DCL algorithms. This is primarily due to the fact that its optimization objective is predominantly driven by position uncertainty. However, for other state components with smaller scales, such as attitude, gyroscope bias, and accelerometer bias, CI-trace performs poorly. The attitude estimation error reaches several tens of degrees and the bias estimates also fail to serve as an inertial alignment correction. This has a detrimental impact on the vehicle: On the one hand, many tasks require not only precise position but also accurate attitude control; on the other hand, inaccurate attitude and bias estimates significantly impair the vehicle’s ability of autonomous navigation. When the vehicle fails to obtain sufficient distance measurements, errors from inertial navigation only quickly accumulate and diverge. CI-det, on the other hand, achieves better estimation results for state components with smaller scales including attitude and IMU biases. However, it performs the worst in position estimation, especially for vehicle 3, where its position RMSE is significantly higher than that of other methods. The difference between vehicle 3 and vehicle 6 also indicates the instability of CI-det in position estimation. In contrast, DCL WCI achieves a balanced and stable estimation across all state components, yielding satisfactory results.
TABLE III: RMSE, RMTE and NEES for Different Methods and State Components (Averaged over all time steps)
Method Position [m] Velocity [m/s] Attitude [deg] Gyro Bias [deg/h] Accel Bias $\boldsymbol{\mu}$g]
RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES
 NCL 4750 3886 25.172 138.0 113.8 56.29 30.97 27.94 4.128 2494 2494 3.000 2424 2425 2.998
CCL EKF 0.078 0.071 3.902 0.035 0.032 5.139 1.020 0.758 5.945 290.2 274.5 5.364 599.6 461.8 11.45
DCL CI-trace 0.305 1.419 0.282 0.369 1.923 0.203 73.876 216.4 3.495 3.4E+04 1.1E+05 1.631 7.6E+04 2.3E+05 1.688
DCL CI-det 0.636 2.305 0.319 0.146 0.894 0.131 2.930 32.72 0.186 521.7 4342 0.183 1307 1.6E+04 0.175
DCL WCI 0.259 1.772 0.196 0.108 0.909 0.084 2.854 41.53 0.081 559.6 7063 0.081 1257 1.8E+04 0.076
(a) RMSE, RMTE and NEES of vehicle 1
Method Position [m] Velocity [m/s] Attitude [deg] Gyro Bias [deg/h] Accel Bias $\boldsymbol{\mu}$g]
RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES
 NCL 5.837 2.316 107.5 0.230 0.146 11.20 1.286 0.934 6.389 317.0 306.1 7.180 874.6 712.3 10.38
CCL EKF 0.074 0.062 3.862 0.028 0.026 4.898 0.886 0.695 5.094 258.4 260.5 4.918 550.6 431.3 9.634
DCL CI-trace 0.217 0.632 2.771 0.190 0.464 2.549 46.88 40.91 26.30 1.7E+04 1.7E+04 7.584 3.2E+04 3.2E+04 20.84
DCL CI-det 2.301 1.915 6.238 0.124 0.129 3.469 1.072 0.947 3.912 297.6 310.1 4.427 743.1 705.6 6.028
DCL WCI 0.201 0.727 1.620 0.053 0.113 1.339 1.187 1.636 1.241 303.0 409.8 1.281 682.4 1116 0.882
(b) RMSE, RMTE and NEES of vehicles 3
Method Position [m] Velocity [m/s] Attitude [deg] Gyro Bias [deg/h] Accel Bias $\boldsymbol{\mu}$g]
RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES RMSE RMTE NEES
 NCL 1.266 1.040 4.584 0.131 0.123 4.035 1.354 0.991 4.821 364.9 317.0 5.553 769.3 729.5 5.502
CCL EKF 0.063 0.059 3.428 0.029 0.026 4.307 0.935 0.680 4.861 270.1 256.3 5.502 498.1 445.0 6.651
DCL CI-trace 0.260 0.481 2.207 0.194 0.341 2.983 26.67 30.39 32.32 8217 8524 5.647 1.3E+04 2.4E+04 4.188
DCL CI-det 0.384 0.556 2.412 0.075 0.094 2.370 1.273 1.075 3.181 347.4 343.6 3.220 660.7 749.6 2.627
DCL WCI 0.203 0.464 1.702 0.066 0.106 1.383 1.394 1.675 1.555 364.7 446.1 1.349 695.3 1144 0.895
(c) RMSE, RMTE and NEES of vehicles 6
To vividly compare the differences in state estimation capabilities between various methods, we use the reciprocal of the averaged RMSE of each state component as a measure of capability. Since the CCL EKF method provides the most accurate state estimation and, therefore, the highest capability, we normalize the capability values using the centralized EKF method as the reference. The resulting radar chart is shown in Fig. 7. From this chart, it is evident that, compared to DCL methods based on traditional CI, DCL WCI achieves a more balanced and effective estimation for all state components, addressing the performance shortcomings of traditional methods in estimating certain states.
Figure 7: The radar chart for state estimation capability of different methods. The capability values of DCL WCI almost entirely encompass those of traditional CI across all five dimensions, effectively addressing the significant shortcomings of the latter. This demonstrates the effective balance that DCL WCI achieves in state estimation.
We also compared the impact of the hyperparameter $T_{a}$ chosen for the weighting matrix in DCL WCI. The results are shown in Fig. 8, where we evaluate the performance of DCL WCI with different weighting matrix by selecting different values of the hyperparameter $T_{a}$ in (30) and (32). Since CI-based methods are conservative estimators, RMTE and RMSE do not perfectly align with each other, and WCI primarily optimizes RMTE. A larger value of $T_{a}$ indicates a greater emphasis on the long-term performance of the inertial navigation system during optimization, thereby assigning more weight to the uncertainty of states with higher-order relationships to propagation errors, such as attitude and IMU biases. This is reflected in the figure, where, as $T_{a}$ increases, the RMTE of position gradually increases, while the RMTEs of attitude and IMU biases decrease.
Figure 8: State estimation performance versus different values of $T_{a}$ for DCL WCI
At the end of this subsection, we evaluate the advantages of the concurrent update strategy in the proposed DCL algorithm. We compare it with the sequential update counterpart, where each vehicle updates its states immediately upon receiving one measurement. As shown in Fig. 9, it is evident that the proposed concurrent update aligns better with the characteristics of CI fusion, exhibiting lower conservativeness and superior estimation performance compared to sequential updates.
Figure 9: Comparison of update schemes: concurrent updates versus sequential updates
V-B Comparison Between DCL and CCL
In this subsection, we compare the DCL and CCL algorithms from two aspects. On one hand, we evaluate their robustness, focusing on the algorithms’ ability to adapt to scenarios involving communication packet loss and changes in swarm topology. This reflects whether the algorithm can remain effective and stable under complex and harsh conditions. On the other hand, we assess the computational and communication overhead of each method, which is a crucial factor influencing the scalability of the swarm and determines whether the algorithm can be applied to large-scale systems.
V-B1 Robustness
Figure 10: Position error comparison between DCL and CCL algorithms. After 30 seconds of motion, there is a 5% probability of data transmission failure from vehicle 1 to the data center. At 90 seconds, vehicle 1 goes offline and loses connection with both the computing center and other vehicles. It then rejoins the network at 110 seconds, restoring its connection with other vehicles.
The CCL algorithm requires all information to be gathered at the computing center. If packet loss or certain vehicle failure occurs, it not only halts the state estimation of the affected vehicle but also leads to incorrect correlation maintenance between vehicles. This, in turn, gradually affects the state estimation of other vehicles as well. We evaluate the robustness of the distributed and centralized algorithms under two conditions: (1) Communication packet loss, where data from certain nodes fails to reach the computing center, and (2) Changes in the network topology, where nodes go offline or join the system. It is assumed that, starting from 30 seconds, there is a 5% probability of data transmission failure from vehicle 1 to the data center. At 90 seconds, vehicle 1 goes offline, losing its connection with the computing center and other vehicles. vehicle 1 then re-joins the network at 110 seconds, restoring its previous connections.
We select vehicle 1 and vehicle 3 as representative examples for analysis. As shown in Fig. 10, we observe that after 30 seconds, not only does the position estimation error of vehicle 1 by CCL EKF method gradually increase, but it also skews the estimation results of other vehicles (such as vehicle 3) that have not experienced packet loss. At this point, the actual RMSE exceeds the estimated RMTE, indicating inconsistent estimates. In contrast, the DCL methods are completely unaffected by this issue.
During the 90-110s period, vehicle 1 is fully offline and disconnected with all anchors and other vehicles, relying solely on its own IMU outputs for state propagation. The position error increases rapidly. However, methods like DCL WCI and DCL CI-det, which achieved balanced state estimation in earlier stages, preserve smaller attitude and IMU biases errors and uncertainties, leading to slower error accumulation during the autonomous navigation phase. Furthermore, when vehicle 1 reconnects, its state is not updated by the centralized algorithm during its offline period. As a result, when vehicle 1 re-joins, its differential distance measurements with other vehicle exhibit significant bias, which in turn skews the position estimates of the other vehicles (such as vehicle 3). In contrast, the DCL algorithm allows the estimation of each vehicle to remain independent, making it more robust.
V-B2 Computation and Communication Overhead
To evaluate computational complexity, we compared the total processing time of the DCL algorithm and the CCL algorithm under different swarm sizes using the same platform (MATLAB R2022b, Intel Core i7-12700, 32GB RAM), as shown in Fig. 11. The total computational cost of the DCL algorithm increases linearly with the number of vehicles. When averaged on each vehicle, the computational overhead for its own state estimation remains almost constant as the swarm size increases, making it more suitable for large-scale swarms. In contrast, as the number of vehicles increases, the dimension of the covariance matrix in the CCL algorithm grows linearly, resulting in a cubic growth in computational cost (since the computational complexity of matrix multiplication and inversion involved in state prediction and update are all cubic in terms of matrix dimensions [35]). As a result, the processing time grows rapidly, making it difficult to scale to large swarms.
As for the communication load, the CCL algorithm requires high-frequency IMU data and measurement data to be collected at the computing center. Additionally, vehicles rely on messages from the computing center to obtain their state estimates. Some distributed implementations of CCL algorithms can also broadcast their state and state transition matrices to track the correlations between vehicles [36], which, although optimized to some extent, still require the transmission of large amounts of data. In contrast, our distributed approach only requires each vehicle to broadcast its position and attitude estimates, along with their variances, amounting to just a few dozen bytes of data, significantly reducing the communication burden of the system. Since this conclusion is evident, no separate simulation experiments are conducted for further analysis.
Figure 11: The total processing time of the DCL algorithm and the CCL algorithm under different swarm sizes
VI Conclusion
In this paper, we propose a WCI-based DCL algorithm designed for MVS operating in 3D scenarios. Unlike existing studies that treat CI merely as a plug-in fusion method, we focus on advancing the CI-based data fusion process for 3D cooperative localization. We find that traditional CI methods fail to achieve satisfactory performance in 3D cooperative localization due to significant scale and observability disparities among state components. To address this, we introduce a weighting mechanism, specifically WCI, to range-based DCL in 3D scenarios. Specifically, we design a multi-measurement fusion strategy based on the characteristic of CI fusion and a weighting matrix selection strategy guided by the INS error propagation principles. Simulation results confirm that the proposed DCL WCI algorithm outperforms DCL methods based on traditional CI in 3D cooperative localization tasks, achieving balanced and improved state estimation across different state components. The proposed concurrent fusion strategy yields higher precision and tighter uncertainty bounds compared to sequential updates. Additionally, the distributed framework demonstrates greater robustness and scalability, making it well-suited for large-scale MVS. Future work will focus on developing dynamic weighting matrix selection strategies to adapt to changes in vehicle motion and cluster configurations, as well as exploring more advanced distributed data fusion methods to improve the conservativeness of CI.